### v1

# Scaling the Authoring of AutoTutors with Large Language Models

Abstract.
Large Language Models (LLMs) have found several use cases in education, ranging from automatic question generation to essay evaluation. In this paper, we explore the potential of using Large Language Models (LLMs) to author Intelligent Tutoring Systems. A common pitfall of LLMs is their straying from desired pedagogical strategies such as leaking the answer to the student, and in general, providing no guarantees. We posit that while LLMs with certain guardrails can take the place of subject experts, the overall pedagogical design still needs to be handcrafted for the best learning results. Based on this principle, we create a sample end-to-end tutoring system named MWPTutor, which uses LLMs to fill in the state space of a pre-defined finite state transducer. This approach retains the structure and the pedagogy of traditional tutoring systems that has been developed over the years by learning scientists but brings in additional flexibility of LLM-based approaches. Through a human evaluation study on two datasets based on math word problems, we show that our hybrid approach achieves a better overall tutoring score than an instructed, but otherwise free-form, GPT-4. MWPTutor is completely modular and opens up the scope for the community to improve its performance by improving individual modules or using different teaching strategies that it can follow.
MWPTutor
I see your solution. You seem to have made some errors. How about we start fresh and do it step by step?
Okay, I’m not sure where to start though.
What should be the first step?
I think Cody’s starting amount of tickets, which is 49, should be the first step.
Tell me this instead: How many tickets does Cody have left after spending 25 on a beanie from his original 49 tickets?
Oh, he would have 24 tickets left.
24 is correct, well done. What is the next step?
Next, he won 6 more tickets, so I think we add that to the 24 he already has.
Okay, can you do that calculation?
Yes, 24 + 6 is 30, so Cody would have 30 tickets in total.
30 is the correct answer to the problem. Well done!!! Don’t hesitate to contack me again if you have any troubles.
GPT4 Tutor
Hi! Could you walk me through your solution?
I’m not sure where to start. I know Cody won 49 tickets and spent 25 on a beanie, but I don’t know how to figure out how many he has left.
That’s a great starting point. So, if Cody had 49 tickets and he spends some on a beanie, what kind of math operation would you use to find out how many tickets he would have left?
Subtraction, I think.
Exactly, subtraction is the right operation. Can you try subtracting the number of tickets Cody spent on the beanie from the number of tickets he initially won? What do you get?
— 12 more tutor utterances —
28 tickets.
That’s correct. Now, if we subtract another ticket from the 28, what number do we get?
27 tickets.
Yes, that’s right. And another ticket taken from 27? What does that leave us with?
26 tickets.
Exactly. And if we take one more away from 26, where does that get us?
25 tickets.
— 13 more tutor utterances —
1. Introduction
Computer-based automated tutoring has long been seen as the method to scale up the benefits of expert human tutoring to millions of students, a problem originally posed by Bloom (Bloom, 1984). Of various attempts to do so, the AutoTutor family (Nye et al., 2014) is perhaps one of the more notable ones. This family of tutors simulates natural language conversations with students based on pre-defined scripts, aiming to emulate one-on-one tutoring. While the AutoTutor family has shown substantial learning gains (Nye et al., 2014), it is difficult to expand it to different topics. Each expansion requires a lot of expert labor to incorporate the required domain knowledge for every concept in the curriculum and appropriate teaching strategies (Koedinger et al., 2004).
Recent developments in Large Language Models (LLMs), such as GPT (OpenAI, 2023), or FlanT5 (Chung et al., 2022), have opened up a new venue for computer-based tutoring where the LLMs play the role of the teacher. While previous work has shown the success of LLMs for tutoring (Rajala et al., 2023; Baidoo-Anu and Ansah, 2023) the non-modular black-box nature of these models makes it impossible to make incremental improvements or control the pedagogy it pursues in assisting the student. The uninterpretable and uncontrollable nature of LLMs makes it difficult for experts or teachers to make changes in their strategy. It is also difficult to develop trust towards such models, not least because they might be open to jailbreak attacks (Liu et al., 2023) as well as privacy concerns. Finally, the proprietary nature of highly performant LLMs means that the tutor needs a persistent internet connection and an OpenAI account to function, and we are highly dependent on OpenAI policies. This could be bypassed by fine-tuning a local open LLM, but doing so leads to either prohibitively high resource requirements, or significantly worse performance (see FlanT5 vs ChatGPT in (Macina et al., 2023)).
The design process of AutoTutor and other similar intelligent tutoring systems can be split into two distinct stages, carried out by different, yet potentially overlapping, sets of stakeholders. The first stage is the general framework level design, carried out mostly by learning scientists (possibly along with computer scientists). This includes designing the pedagogy, the mechanisms required to interact with the student, and potentially even authoring tools. This is typically done only once for an entire domain. The second stage, performed by either teachers or domain experts, is to author the scripts111see Appendix A in (Graesser et al., 2000) for an example script, which needs to be done separately for every topic, question, or concept. In the standard framework, both these stages are done by humans. In contrast, in the LLM-only framework discussed in the previous paragraph, both these stages are carried out (implicitly or explicitly) by LLMs.
In this paper, we consider a mixed approach where we only delegate the job of domain experts to an LLM, while the first stage is still left to learning experts. This idea has previously been hinted at by Cai et al. (Cai et al., 2019), who propose using a question-answering system to automatically list potential misconceptions for a given topic. We extend this idea to have an LLM author the entire script for the tutoring system, which is accurate enough to be deployed without human validation but is fully open to human validation.
Our contributions in this paper are as follows:
We adapt the Dialog Based tutoring framework used by ITS’s to make it possible for a Large Language Model to author it.
We design MWPTutor, a tutor for math word problems based on the above principles.
We compare the performance of MWPTutor with that of GPT4 being used as a tutor both quantitatively and qualitatively.
2. Background
Tutoring is an integral part of teaching where the teacher attempts to identify confusions or misconceptions a student might have, and rectify them. This is done in a way to minimize their future re-occurrence. Bloom et al. (Bloom, 1984) showed that tutoring is extremely important for education, as a student receiving one-on-one tutoring can perform much better than a student only getting lectures. As opposed to lecturing, tutoring is more personalized, and therefore only feasible in smaller groups. This lower teacher-to-student ratio needs to be balanced by a higher number of tutors, which is expensive and not always feasible.
2.1. Rule-Based Systems
To combat the above scarcity, Intelligent Tutoring Systems (ITSs) were already being developed since the early 1980s (Sleeman and Brown, 1982). The idea was to roleplay a human tutor with a computer program that could interact with the student and fulfill the requirement of personalized tutoring. The computer program may not fully replace human tutors but often provides an affordable and scalable addition. An important aspect that differentiates ITSs from simply showing the students the problem solution is their ability to provide step-by-step feedback which promotes long-term learning as opposed to short-term problem-solving success. This was often implemented using rules that check for certain errors using string-matching and act on them with pre-defined templates. For example, LISP Tutor (Anderson et al., 1985) contains rules to account for common programming errors made by students.
With advancements in AI, ITSs transitioned to dialog as a method of tutoring, moving ITSs from being information-delivery systems to dialog-based agents. One of the key players in this field was the first version of AutoTutor (Graesser et al., 2001), which replaced the rules in traditional ITSs with a full finite state transducer which they called the Dialog Advancement Network. This allows the students to interact in natural language and the tutor processes their inputs using methods like latent semantic analysis and attempts to respond using a set of predefined rules. Over the next two decades, various tutors such as the AutoTutor explored different domains, tutoring strategies, affective statements, etc. to become one of the most comprehensive ITSs (see (Nye et al., 2014) for the summary of developments).
With concepts and tutoring strategies getting more complex, writing the rules governing the finite state transducers quickly becomes time-consuming. Not knowing the student misconceptions in advance, domain experts have to come up with rules for numerous possibilities, introducing a different set of scaling issues. One hour of interaction with a student costs over 100 hours of domain expert time to author the ITS (Koedinger et al., 2004). This is in addition to the time spent by learning scientists designing the pedagogy and dialog strategies. However, this is usually done only once for the topics in a subject area. As a result, this limits the AutoTutor to a small set of subject areas, restricting their impact. Nevertheless, there are many works in the space of rule-based ITS (Paladines and Ramirez, 2020), which have been reasonably successful in their domains.
2.2. Large Language Models
Large Language Models (LLMs) are Transformer-based models that serve as general-purpose tools in natural language processing (Bommasani et al., 2022). With recent improvements in LLMs and the availability of LLM-as-a-service models like GPT (OpenAI, 2023) and BARD (Yarkin and Coon, 2022), it has become easy to use LLMs as tutors by prompting them with the desired qualities of a tutor and then letting them interact with a student. Some works in the programming domain like GPTutor (Chen et al., 2023) and TestGPT (Phung et al., 2023) attempt to use GPT4 similar to information delivery systems. There have also been systems designed to function as Dialog agents, like Ruffle&Riley (Schmucker et al., 2023) which includes multiple GPT4 agents functioning with different prompts, and SPOCK (Sonkar et al., 2023) which attempts to distill out tutoring knowledge from GPT4 into a smaller LLM. A notable commercial usage of a LLM Tutor is seen in Khanmingo,222khanacademy.org/khan-labs which uses the GPT4 backend for an interactive tutor.
While LLM tutors solve the scaling issues, it introduces a whole new set of complications, primarily related to its probabilistic output and proprietary models. The behavior of GPT4 can only be controlled by making changes to its initial prompt or by inserting prompts in between, making it hard to introduce the nuanced pedagogical strategies developed by learning scientists over the last two decades. Even for simpler strategies, its probabilistic nature leaves it open to being exploited by users who behave in unwanted ways, often referred to as jailbreak attacks (Liu et al., 2023). Even for benign users, GPT4 can often hallucinate incorrect information and output it with confidence (Zhang et al., 2023; Cui et al., 2023) which may adversely affect the student. All this, in addition to their hard-to-understand nature, leads to a distrust of GPT4 among educators and policymakers, causing it to be banned in several US public schools.333bestcolleges.com/news/schools-colleges-banned-chat-gpt-similar-ai-tools
2.3. Our work
The main benefit of the fully human-authored setting comes from the fact that it is easy to control, easy to interpret, and therefore easy to trust. Most of these benefits come from the pedagogy design stage, and not from the work of the domain experts. The LLM-based approaches, on the other hand, mostly speed up the domain experts’ tasks. To bring together the best of both worlds, we propose an expert-designed pedagogy to be assisted by an LLM in the dialogue part. The rest of this paper is structured as follows: Section 3 introduces our general design which applies to several domains, Section 4 describes the specific design of MWPTutor, Section 5 quantitatively evaluates MWPTutor, and Section 6 looks qualitatively at weaknesses of both GPT4 and MWPTutor.
3. Design Principles
While this work mainly focuses on Math Word Problems tutoring, our broader goal is to present a paradigm that can be used across different subjects. Before the math-specific aspects, we first describe the model design principles. The output is the implementation of a pedagogical goal as a template state space for a finite state transducer similar to dialog advancement networks, which can then be filled by an LLM. For the pedagogical strategy, a good doctrine to follow can be Get the student to the correct answer. If possible, do so without telling them the answer outright
A finite state transducer is characterized by the combination of state space and a transition function. The state space can be seen as the memory of the finite state transducer, which stores one or more variables, each of which can take two or more values. In our case, we have two variables, the solution step and the strategy, which factorizes our state space into two subspaces, detailed in Section 3.1 and Section 3.2 respectively. The transition function specifies the next stage given the current state and the (student) input and is highly domain-dependent, but we provide examples as they come up alongside the state space descriptions
First, consider a math word problem and a history question to illustrate the possible range of topics which a tutor might apply.
The optimal tutoring support differs between the two examples. While the math word problem starts with subquestions to compute intermediate quantities, answering the history questions first requires providing historical context.
3.1. Solution Decomposition
To begin creating the state space for a finite state transducer, we first need to list the solutions for the problem, creating the solution step space. As a first design principle, we want to have the solution broken down into multiple parts. For the math word problem example, this could look like a tree:
To complete the question, the student needs to perform A1 followed by A2. Alternatively, they can complete B1 and B2 in any order and then B3. We transition out of a step if the student arrives at the right-hand side of the equation in the (sub)question. These dependency structures could look very different based on the domain, like for the history question, for example:
In this case, a correct answer requires the student to complete any 4 out of C1-C6 in any particular order to be considered successful. The transition function here is more complicated, but one way would be to check if all underlined key terms have been mentioned.
Splitting a problem into smaller parts is beneficial for both human learning (Rich et al., 2018; Song and Becker, 2014) and for LLM solvers (Dhuliawala et al., 2023; Nye et al., 2021). Getting an LLM to split the problem into smaller parts allows us to thereafter focus the LLM on smaller parts of the solution, which can also reduce context lengths. In AutoTutor terminology, these solution steps would correspond to “expectations”.
3.2. Pedagogical Strategy
Having broken down a problem into steps, we now need to have a strategy to help the student achieve each step. The general idea is to start vague, hoping for the student to do as much of the work as possible, but then keep getting more specific until the student can complete the step. For the math word problem example, a toy strategy could be:
This set of pedagogical strategies forms our strategy space. We can transition from D1 to D2 and then to D3 if the conditions for transitioning to the next solution step are not met, at which point we loop in D3. For the history example, we have:
The actual set of moves can be quite complex. For P2, we would need separate utterances for every possible combination of missing keywords. Ideally, these moves should be designed by learning scientists as was the case for AutoTutor.
The overall state space is the cross-product of the solution state space and the strategy space. Figure 2 shows what the full state space of our MWP example could look like. For each element of the strategy space, we would require prompt templates that can be filled with the solution step and passed on to LLMs to generate the actual utterance. Running each prompt for every step in the solution space maps every element in the state space to an utterance. Not all utterances need to be unique, and some, like P1 above, might not necessitate an LLM to be generated.
3.3. Scaffolding the LLM
While modern LLMs fine-tuned with RLHF are quite good at following instructions, they can at times struggle with more complex tasks. Because of the probabilistic nature of LLMs, reattempting the same problem can lead to different answers. To take the best advantage of these facts, we want to have automated checks for every LLM turn in the strategy space, so calls can be repeated until some checks have been met. For example, D3 could require that the numerical answer (450 for A1) is included in the utterance, while P2 could require the missing keywords to be absent. These checks can take different forms depending on the domain, like limiting the length of an utterance or avoiding the use of more complex language.
3.4. Control and Reliability
One of the primary goals of having LLMs author a state space rather than having them interact directly with the students is that in this way, one can examine and trust the model without having to necessarily trust the LLM used to generate it. Based on this, we store and make it possible to inspect beforehand the entire utterance space. Changes to the utterances themselves can be made on-the-go, e.g. if a teacher feels it is better to split our example in P2 into two sentences, they can simply edit it to say “Who was assassinated?” and “Who carried out the assassination” with no further changes required. Changes to the solution space, however, need the utterances to be regenerated.
Caching the utterance LLMs like GPT also makes the system immune to changes made to GPT in terms of availability, cost, or rate limits. For open LLMs, this allows for the use of Larger LLMs which not every user can run locally. The most important benefit, however, is the potential for teachers/domain experts to edit the state space should they want, therefore allowing us to combine human expertise with machine efficiency. Live calls to an LLM can often help and can be incorporated into the state space. For example, in the MWP we could make a call to GPT to take over a step in between D2 and D3 if we want to minimize the chance of D3. Should live calls be implemented, however, there should ideally be non-LLM fallbacks ie we should have cached and/or procedural alternatives available should a call to an LLM not be possible at deployment time.
4. MWPTutor
We now detail the design process of our system MWPTutor, an instantiation of the above principles. MWPTutor is adapted to help students solve math word problems. The scenario, originally defined by MathDial (Macina et al., 2023), involves a student approaching a tutor after unsuccessfully attempting to solve a math word problem. The student’s incorrect solution is provided as a part of the input. GSM8k (Cobbe et al., 2021), the source for MathDial’s problems includes a single step-by-step solution to each problem, which can be seen as a fulfillment of the solution step space. However, to not become constrained to the availability of such solutions, and to allow diverse solution paths, we restrict MWPTutor to only the final numerical answer to the problem, and generate the solutions ourselves.
symbol indicates student utterance inputs, while the
symbol indicates model output. The solution step space is collapsed for clarity
4.1. Solution Decomposition
We instantiate the solution steps as short pieces of text describing a step in the math word problem, including a single mathematical calculation. We generate multiple solution paths with our backend LLM, with each step needing to contain a single calculation demarcated in double angular braces (e.g. , see Table 3). This requirement standardizes the information density of a step and also keeps things consistent with GSM8K which uses a similar notation. It also allows us to check for calculation errors and discard erroneous paths. Once multiple solution paths have been obtained, we discard the paths not leading to the final correct solution, and group the rest together using exact string match on the calculations into a “tree” of solution steps. Our running example, except for the angular brackets, fits this criterion as a tree with two branches at the root and no branches elsewhere.
4.2. Solution Alignment
Although not a part of design principles, we wish to not make the students repeat a step they have already performed correctly. To achieve this, we match steps in the original student solution with the closest path in the solution step space. While AutoTutor achieves this using latent semantic analysis and string-matching techniques, we found that modern counterparts, such as BERTScore (Zhang et al., 2020) are surprisingly not an apt replacement, perhaps because GPT4 outputs very long step descriptions, and BERT does not weigh and compare numbers and mathematical operations well.
MWPTutor currently assumes that the student will present their solution with one step in each line. For each student step, we check if the right-hand side of any of the possible follow-up steps is present in it. If more than one follow-up matches, we pick the one with the highest BERTScore. We then proceed down that node of the solution tree and the next student step. If no follow-up matches, the current step is considered to be the first point of error. The pseudo-code for this is shown in Algorithm 1.
Input: gold and student solutions.
Output: aligned solution step, if exists
Filter out next solution steps whose RHS is in the current student step
Move to next step
4.3. Pedagogical Strategy
To instantiate our pedagogy, we take inspiration from AutoTutor, where the primary sequence of steps is a pump→hint→prompt→ assertion (Section 2 of (Graesser, 2016)). Graesser et al. in their work Teaching Tactics in Autotutor (Graesser et al., 2000) define hints as “questions that lead the student to the desired information”, and assertions (which they call elaborations in this work) as “succinctly assert the desired information.” For prompts, they use fill-in-the-blanks type utterances leaving out a keyword. In our case, the “keyword” is the equation, so instead we explain the calculation to be performed in a natural language. These are preceded by a “pump”, which is the generic utterance “What should be the next step?” The overall strategy then proceeds as follows:
| Pump |
asks the student to propose the next step. If the student proposes the step correctly, we ask them to solve it. |
| Hint |
is only triggered if the student fails to propose a valid next step, or fails to execute the step they propose. It essentially gives out the step and asks the student for the calculation. This utterance should take the form of a question, and the only way to correctly answer the question should be to successfully execute the step. |
| Prompt |
is triggered if the student fails to execute the step after the hint. This should push the student closer to the answer, without actually revealing the answer. |
| Assertion |
is triggered if the student fails to answer even after the prompt. This should simply give out the answer, and have the student acknowledge it. |
In our running example, hints, prompts, and assertions would correspond to D1, D2, and D3 respectively. The pump is not included.
For scaffolding the LLM, both hints and prompts are required to avoid explicitly outputting the right-hand side of the equation in the step, with the hint additionally having to avoid saying any words that indicate arithmetic operators, currently “add, subtract, sum, multiply, product, divide, quotient.” The assertion, on the other hand, must contain the right-hand side. For pumps, there is no use of an LLM, so we do not need a check.
4.4. Answer and Path Detection
To execute the pedagogical strategies described above, we need to be able to detect if the student has given a correct answer. Further, the hint state should only trigger if the student fails to propose a correct follow-up. For this, we need to have a way to match paths. We run into the same matching problems described in Section 4.2, and resort to similar basic techniques. The correct answer detection algorithm is shown in Algorithm 2. For states other than pump, it takes a single reference step and checks if its RHS is present in the student utterance. For the pump state, all possible followups of the last step are passed as reference steps. If it can find the RHS of any of these in the student utterance, it marks that as the next step and proceeds to the next step. If not, it considers a path match to have occurred if every operand on the left-hand side of some step is present. If multiple steps fulfill this, or no step does, we pick the one with the best BertScore alignment to the student utterance.
The fact that we currently ignore all non-numeric content in the student response is one of the major limitations in our work, as it can miss a large number of “correct” utterances from the student, but we defer the resolution of this to future work.
Inputs: CurrentSteps from Solution step space, and Student Utterance
Output: Degree of the match, Best matching step
CurrentSteps is a singleton except for “pump.”
Go to next solution step
4.5. Putting it All Together
Figure 3 summarises the overall state space of MWPTutor. It can be seen as a simplified version of the state space of AutoTutor (see Figure 1 in (Cai et al., 2019)). The expectation matching is implicitly carried out by the solution aligner and the pump state. To obtain to full finite state transducer structure, one must roll out the states along the solution space. Excluding the unlikely case where the student fails to repeat the answer after the assertion state, the overall conversation should not last more than utterance pairs, where is the number of steps.
5. Experiments
In this section, we the performance of MWPTutor and GPT4 in terms of how well they can follow our previously mentioned doctrine. We test both models in the interactive setting proposed by Macina et al. (Macina et al., 2023) which allows us to safely compare imperfect models without having a negative impact on human users.
5.1. Setup
5.1.1. Setting.
The interactive tutoring setting was proposed in MathDial as a faster, cheaper, and safer alternative to human experiments, where instead of interacting with human students, a Tutoring System interacts with another LLM role-playing a student. In their data collection experiments, MathDial found that most teachers believed that the student LLM behaved very much like an actual student in most cases. We adopt their setting as is for our experiment section.
Note that for an LLM, the interactive tutoring setting is somewhat related to the MWP solver setting often used as a metric to measure the reasoning abilities of said LLM, but there are significant differences. Unlike the MWP solver setting, the final answer is already provided, so the LLM does not need to come up with it. The difficulty arises from the fact that this LLM needs to get the student LLM to get to this answer, and the student LLM is prompted to not budge from its starting wrong answer without good reason. Despite these differences, some skills like Math Reasoning Abilities transfer between the two settings, so we can expect the final performances in both settings to correlate with each other.
5.1.2. Models.
At the time of testing, GPT4555gpt-4-1106-preview queried in January 2024 was considered to be one of the best-performing LLMs, so we use it both as a backend for MWPTutor and as the baseline single prompt LLM. To best align the single prompt GPT4 with the MWPTutor doctrine, we ask it to roleplay a tutor, setting the primary objective of getting to the correct answer as the goal of the conversation, and the secondary objective of not revealing the answer as a personality trait. Further, to test the potential of an open-source LLM as a backend, we have a second version of MWPTutor with a MetaMath-13B (Yu et al., 2023). MetaMath is a version of Meta AI’s open source LLM LLaMa666https://llama.meta.com/ fine-tuned to solve math word problems, and is, to the best of our knowledge, the best open LLM for solving math word problems. We refer to these three models as MWPTutor, GPT4, and MWPTutor respectively.
5.1.3. Datasets.
To implement MWPTutor, we require a dataset with math word problems and their solutions, as well as incorrect student solutions. We make use of the test sets of two datasets, namely MathDial (Macina et al., 2023) and MultiArith. MultiArith consists of simple Math Word Problems that are usually solvable in 1 to 3 steps. GPT3 was already able to solve over 90% of these problems with the right prompts (Kojima et al., 2023), and in our tests, we found that GPT4 was able to get a perfect score in the MWP solver setting. With this dataset, we want to test if we can achieve the same performance in the interactive tutoring setting. This dataset does not provide any wrong solutions needed to instantiate the student model, so we make use of errors made by a MetaMath solver as the original student solution where possible, and leave the remaining questions as “No idea how to begin”.
MathDial consists of problems from the pre-existing GSM8k (Cobbe et al., 2021). Problems in GSM8k are meant to be solvable by a bright middle school student using only basic arithmetic operators (, , , ), taking up to steps to solve. Of these, MathDial picks problems that have official solutions taking steps or fewer, that GPT3.5 consistently failed to solve. This makes MathDial a significantly harder dataset to teach, and results on this dataset are expected to show the differences between the models more clearly.
5.1.4. Metrics.
Keeping in line with our doctrine, our primary metric of evaluation is the Success Rate (S), which is measured as the fraction of conversations where the student says the correct answer eventually. The secondary metric is Telling Rate (T), which is the fraction of conversations where the tutor did not reveal the answer. The tutor is said to have revealed the answer when they say the final answer before the student has said it, and the student succeeds in getting to the answer thereafter. Both these metrics were originally defined by Macina et al. (Macina et al., 2023), but they do not offer a way of combining them into a single score. It is quite easy to increase success at the cost of increased telling. Therefore we also calculate the Adjusted Success (S-T) rate as the fraction where the student succeeded without having been told the answer and use the harmonic mean between this number and the success rate as an overall Tutoring Score (). We also offer utterance level evaluations in Section 5.3, but since we are unaware of the utterance level strategies from GPT4, we do it only for our model, and the metrics for the same are defined later.
5.2. Conversation Evaluation
We evaluate both success and telling for the MultiArith dataset by directly looking for the answer in student and tutor utterances respectively, and calculate the overall scores based on that. The top half of Table 1 summarises our results. All models are able to transfer their near-perfect solving scores to near-perfect tutoring scores. GPT4 performs slightly worse, although it has a telling rate. Both versions of MWPTutor have a perfect success rate. It must be noted that the metric is noisy since it only requires the number to match, with no restriction on the context, e.g. in the problem “How many groups of can be made out of boys and girls?” if the tutor mentions there are boys, we would count it as telling as the final answer happens to be . This means that the real rates could be somewhat lower for all models, and the difference between them is not significant. Nevertheless, such errors are unlikely to be frequent, and all models work reasonably well on these problems.
| MWPTutor | GPT4 | |||
| Metric | LLaMa | GPT | ||
MultiArith |
Success Rate | 1.00 | 1.00 | 0.98 |
| Telling Rate | 0.01 | 0.01 | 0.00 | |
| Adjusted Success | 0.99 | 0.99 | 0.98 | |
| Tutoring Score | 0.99 | 0.99 | 0.98 | |
MathDial |
Success Rate | 0.91 | 0.96 | 0.82 |
| Telling Rate | 0.33 | 0.11 | 0.00 | |
| Adjusted Success | 0.58 | 0.84 | 0.82 | |
| Tutoring Score | 0.71 | 0.90 | 0.82 | |
For the harder MathDial dataset, we wanted to avoid matching errors as described earlier, and so have human annotators mark out utterances where the student or the tutor says the final answer. Each conversation was annotated by annotators, and we took the majority in cases where they disagreed.
Table 1 summarises our result. While GPT4 maintains its 0% telling rate, MWPTutor gets more problems correct overall, and more when adjusted for telling. This gives MWPTutor an advantage on GPT4 in terms of the final score. MWPTutor also gets a higher success rate than GPT4, which is somewhat surprising since its backend model is significantly weaker than GPT4 in the MWP Solver setting (Yu et al., 2023). This increased success rate however comes at a cost of a very high telling rate, giving it a significantly poorer tutoring score overall, so we do not recommend the use of MetaMath as the backend model unless GPT4-like models need to be avoided for some external reason. One thing worth mentioning here is that once problematic questions have been identified for either version of MWPTutor, one can have human annotators redo only those questions, thereby increasing the success rate, something that is not possible for GPT4.
5.3. Utterance Evaluation
A major shortcoming of AI-on-AI-tutoring is that the AI student may not be affected by the sometimes subpar tutor responses. This is more common in MWPTutor than in GPT4, as the former can only pick from a pre-defined set of responses and not create new utterances on the fly in response to finer details in the student requirements. To understand if this is a major issue, we hired human evaluators who worked in the education sector and asked them to evaluate utterances made by MWPTutor for MathDial. Because the pumps are generic and very similar to each other and assertions essentially boil down to reading out the answers, we focus only on the hints and prompts in our evaluation.
Across the conversations, MWPTutor and the student model went over 135 solution steps. Of these went to the hint state, while reached the prompt state. We presented each of these utterances to human evaluators, with the past conversation, the original math word problem, and the current solution step as context. We boil down the requirements set on hints and prompts set in Section 4.3 to the following binary criteria which we asked the evaluators to evaluate the utterances on:
Grammatical: The utterance is linguistically and grammatically correct.
Not Reveal: The utterance does not reveal the final answer for the step.
Relevant (hints only): The hint clearly addresses the step, ie, the only way to correctly answer the question is to complete the current step.
Not Redundant (hints only): The hint is not redundant to what the student already said in the previous step ie it makes sense to ask this question after the previous response.
Specific (prompts only): The prompt is specific to the step, and not just a vague lead-on statement.
Helpful (prompts only): The prompt can potentially help the student make progress in the current step, given the previous student’s utterance.
The annotators were paid the Prolific suggested wage of 10 GBP/hour. We show the annotation user interface in Figure 6 and some example annotations in Table 2 and Table 3. Figures 4 and 5 summarise the responses from the evaluators. Almost all utterances are considered Grammatical by the evaluators, which is not surprising given they were generated by GPT4. Most of the utterances are also marked to be non-revealing, which is also not surprising given that any proposed utterances that directly revealed the answer were assigned to be regenerated. The evaluators further seem to agree that most hints were relevant to the step and that most prompts are specific, implying that they do what is appropriate given the current solution step.
The biggest potential weakness of MWPTutor comes in the form of its utterances making sense in the context of previous student utterances due to issues discussed in Section 4.4, which is where the final two metrics come in. In this regard, all prompts are considered to be helpful by a majority of evaluators which is a great sign. Hints, however, are found to be redundant in almost a third of cases, showing that handling of student’s responses to pumps could be better. The current heuristic of matching the left-hand side is too demanding, and highly likely to ignore the student even if they propose a valid next step to the problem, which can break student immersion. Note that of the steps did not proceed to the hint state, implying more than half of the correct proposals by the student are being identified as such. Further, in a vast majority of cases where a correct proposal is misidentified, there is disagreement between human evaluators indicating there might be some merit in the rejection, though this is also the case for a majority of correct rejections.
Overall these evaluations show that utterances generated for MWPTutor by the GPT4 backend are consistent with the strategy definitions, and the transition function uses them reasonably well.
6. Qualitative Assessment of Weaknesses
While both GPT4 and MWPTutor do quite well on quantitative metrics as seen above, these metrics do not fully grasp what it takes to make a good tutor. To better understand the preparedness of our models, we went over all conversations that happened on the MathDial dataset, for both GPT4 and MWPTutor and summarize our findings here.
Firstly, MWPTutor tends to sound monotonous and dry. Part of this is because of the pre-set format for utterances like all hints are hardcoded to begin with “Tell me this instead:,” but this can easily be resolved by having it sample from several rephrasings of the same utterance. More difficult to fix are the repetitive strategy and the lack of finer feedback or any affective statements, both of which are there by design. Introducing finer feedback requires a more fine-grained analysis of student utterances, which we found off-the-shelf NLP methods to be incapable of doing while maintaining a high degree of accuracy. Affective statements, like encouragements, challenges, etc. would further require accurate judgment of student confidence and correctness. In both these cases, we made a design choice to avoid the additional module, as they would create major failure nodes. GPT4 on the other hand, sounds a lot more natural, including finer feedback and probing, and even affective statements, although they are exclusively positive.
Further, since it matches correct answers with limited attention to context, MWPTutor can sometimes get false positives, and inadvertently reveal the answer. Consider the following excerpt:
Here, the student gives the wrong answer of , but MWPTutor mistakes the intermediate calculation of as the correct answer, and effectively reveals it to the student. GPT4 is able to understand context usually, although it is far from perfect. As we discuss later, GPT4 fails much worse on this question too.
While GPT4 conversations flow a lot better and sound human-like, it has a different set of issues. For one, GPT4 tends to overcomplicate questions or make redundant calculations. In one question, where the student failed to convert seconds into minutes, GPT4 gets the student to convert the seconds to hours first, and then back to minutes. This introduced precision errors, which it had to fix by rounding the final answer. In another question GPT4 has a student recalculate a value she had already calculated in a previous step. In yet another question GPT4 makes the student calculate the weekly rate of weight loss, when a simple multiplication of time period by 4 would have yielded the answer. The conversation in Figure 1 is also an example of where this happens in the simpler MultiArith dataset. Since it explores multiple paths and is also limited in the steps it can take, MWPTutor is less likely to run into this issue, and handles the first two cases mentioned here pretty well, though it makes the same error in the last case.
The biggest problem with GPT4, however, continues to be that GPT4 can lead the student to a wrong answer, sometimes even leading them away from the right answer. In all cases where GPT4 fails to get the student to the correct answer, it outright accepts a wrong answer as correct. Of MWPTutor’s failures one was a result of inadvertently revealing the final answer, while the other seems to have been marked as incorrect by the annotators because the tutor resorts to revealing rather quickly. In a further conversation, including the one regarding chicken feed described earlier, GPT4 misdirects the student onto the wrong path but somehow manages to get them to say the correct answer. Finally, in one question, GPT4 rejects a correct equation, and then goes off in a tangential direction, until the maximum limit of utterances is reached (this being the only instance where the utterance limit comes into play). It must be reiterated here that the final correct answer is given to GPT4 as part of its prompt. MWPTutor by design will never okay a wrong answer. The only place where it does go wrong, it accepts an answer of instead of , but this can be chalked down to the fact that the question was asking for the “difference” between two quantities, and did not make it clear whether the sign was required (the first quantity was smaller. Also GPT4 makes the same error. Both are marked as correct by annotators). It must, however, be noted that since it also relies on GPT4 for solution generation, MWPTutor can hallucinate, even if it did not happen in any of these conversations to the best of the authors’ judgement (We attempted to have the evaluators annotate the existence of wrong methods, but found too much disagreement in their responses to consider them viable). If this happens, however, one can fix them by changing the tree the first time they get noticed, something that is not possible with GPT4.
Overall, we conclude that despite good showing on metrics, both methods are still lacking in quality, and decide to not test them on real students until we can significantly improve them.
7. Discussion and Future work
We have stated throughout the paper that although self-sufficient, MWPTutor is still a prototype, and we explained in Section 6 that we still have a long way to go before making MWPTutor viable to be used on real students. In this section, we outline some directions for future work that could bring us closer to these goals.
7.1. Improving the Solution Step Space
The current solution step space takes the shape of a tree where solutions can only diverge from each step. But this is non-realistic as we know that all paths must lead to the same final answer and hence need to converge at some point. Further, the order of carrying out certain steps could be a partial order rather than a total order. This would currently be represented by listing out all possible orders as separate paths, but this causes a blowup in the number of steps. Adding a richer structure to the Solution graph could vastly reduce the size of the solution space.
In addition to this, each step of the solution should be made to follow a more standardized format. Having a more precise definition of the step would allow us to use LSA-like techniques to improve our correct answer detection. Pre-defining possible errors for each step would allow us to prepare to handle said errors. We can also have descriptions of each quantity in the LHS to give more pointed hints.
7.2. Improving the Strategy Space
The current strategy space is linear and does not allow for very targeted reactions to the student utterances or the original student solution. Once we have a better solution step space, we can look for different types of errors and diverge into different Prompts depending on the exact error should the student fail to get the answer correct after the Hint state. There can even be more than one Prompt revealing different amounts of information based on the complexity of the current step. Another thing entirely missing is the “Problematization” strategies described by Reiser (Reiser, 2004) and also used by MathDial. Alongside this, the matching requirements for solution alignment can be made stricter to better know where to start applying our intervention. This could also come with relaxing the format requirement of “one step per line” on the student solution. Both these would shorten the average lengths of the conversation to balance more steps in the strategy space.
Alternatively, matching can be done using smaller specialized LLMs trained for this task, and this would also apply to correct answer detection and path matching
8. Conclusion
In this paper, we propose using LLMs to take over the authoring of the state space for an AutoTutor-like model to minimize human time requirements and build a sample system to showcase our proposal. We find that our model performs somewhat better than a Single Prompt GPT4 model in terms of automatic metrics while maintaining full interpretability and modularity in its state space which GPT4 lacks. Our model also holds up to the desired qualities in its utterances in human evaluation, showing there is high potential in this direction. But at the same time, it still faces issues that become more apparent when going through conversations manually, indicating that there is a lot of work still to be done.
These issues can be fixed by better LLMs and more sophisticated techniques outlined in future work. Several potential steps have already been outlined in Section 4, and we hope to soon have better versions of MWPTutor and other similar systems that can realize one-on-one tutoring at scale.
Acknowledgements.
This work was funded by the Swiss National Science Foundation (Project No. 197155), and a Responsible AI grant by the Haslerstiftung. We extend our gratitude to Jakub Macina from the ETH AI Center, Nico Daheim from UKP Lab, TU Darmstadt, and Farquet Grégoire from Department of Architecture, ETH Zürich for their invaluable feedback.References
- (1)
- Anderson et al. (1985) John R. Anderson, C. Franklin Boyle, and Brian J. Reiser. 1985. Intelligent Tutoring Systems. Science 228, 4698 (1985), 456–462. https://doi.org/10.1126/science.228.4698.456 arXiv:https://www.science.org/doi/pdf/10.1126/science.228.4698.456
- Baidoo-Anu and Ansah (2023) David Baidoo-Anu and Leticia Owusu Ansah. 2023. Education in the era of generative artificial intelligence (AI): Understanding the potential benefits of ChatGPT in promoting teaching and learning. Journal of AI 7, 1 (2023), 52–62. https://dergipark.org.tr/en/pub/jai/issue/77844/1337500
- Bloom (1984) Benjamin S Bloom. 1984. The 2 sigma problem: The search for methods of group instruction as effective as one-to-one tutoring. Educational researcher 13, 6 (1984), 4–16.
- Bommasani et al. (2022) Rishi Bommasani, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S. Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, Shyamal Buch, Dallas Card, Rodrigo Castellon, Niladri Chatterji, Annie Chen, Kathleen Creel, Jared Quincy Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren Gillespie, Karan Goel, Noah Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E. Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, Omar Khattab, Pang Wei Koh, Mark Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D. Manning, Suvir Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Ben Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, Julian Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Rob Reich, Hongyu Ren, Frieda Rong, Yusuf Roohani, Camilo Ruiz, Jack Ryan, Christopher Ré, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishnan Srinivasan, Alex Tamkin, Rohan Taori, Armin W. Thomas, Florian Tramèr, Rose E. Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang. 2022. On the Opportunities and Risks of Foundation Models. arXiv:2108.07258 [cs.LG]
- Cai et al. (2019) Zhiqiang Cai, Xiangen Hu, and Arthur C. Graesser. 2019. Authoring Conversational Intelligent Tutoring Systems. In Adaptive Instructional Systems, Robert A. Sottilare and Jessica Schwarz (Eds.). Springer International Publishing, Cham, 593–603.
- Chen et al. (2023) Eason Chen, Ray Huang, Han-Shin Chen, Yuen-Hsien Tseng, and Liang-Yi Li. 2023. GPTutor: A ChatGPT-Powered Programming Tool for Code Explanation. In Artificial Intelligence in Education. Posters and Late Breaking Results, Workshops and Tutorials, Industry and Innovation Tracks, Practitioners, Doctoral Consortium and Blue Sky - 24th International Conference, AIED 2023, Tokyo, Japan, July 3-7, 2023, Proceedings (Communications in Computer and Information Science, Vol. 1831), Ning Wang, Genaro Rebolledo-Mendez, Vania Dimitrova, Noboru Matsuda, and Olga C. Santos (Eds.). Springer, 321–327. https://doi.org/10.1007/978-3-031-36336-8_50
- Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, Albert Webson, Shixiang Shane Gu, Zhuyun Dai, Mirac Suzgun, Xinyun Chen, Aakanksha Chowdhery, Alex Castro-Ros, Marie Pellat, Kevin Robinson, Dasha Valter, Sharan Narang, Gaurav Mishra, Adams Yu, Vincent Zhao, Yanping Huang, Andrew Dai, Hongkun Yu, Slav Petrov, Ed H. Chi, Jeff Dean, Jacob Devlin, Adam Roberts, Denny Zhou, Quoc V. Le, and Jason Wei. 2022. Scaling Instruction-Finetuned Language Models. arXiv:2210.11416 [cs.LG]
- Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training Verifiers to Solve Math Word Problems. arXiv preprint arXiv:2110.14168 (2021).
- Cui et al. (2023) Chenhang Cui, Yiyang Zhou, Xinyu Yang, Shirley Wu, Linjun Zhang, James Zou, and Huaxiu Yao. 2023. Holistic Analysis of Hallucination in GPT-4V(ision): Bias and Interference Challenges. arXiv:2311.03287 [cs.LG]
- Dhuliawala et al. (2023) Shehzaad Dhuliawala, Mojtaba Komeili, Jing Xu, Roberta Raileanu, Xian Li, Asli Celikyilmaz, and Jason Weston. 2023. Chain-of-Verification Reduces Hallucination in Large Language Models. arXiv:2309.11495 [cs.CL]
- Graesser et al. (2000) Arthur Graesser, Natalie Person, and Derek Harter. 2000. Teaching Tactics in AutoTutor. (01 2000).
- Graesser (2016) Arthur C. Graesser. 2016. Conversations with AutoTutor Help Students Learn. International Journal of Artificial Intelligence in Education 26, 1 (01 Mar 2016), 124–132. https://doi.org/10.1007/s40593-015-0086-4
- Graesser et al. (2001) Arthur C. Graesser, Kurt VanLehn, Carolyn P. Rose, Pamela W. Jordan, and Derek Harter. 2001. Intelligent Tutoring Systems with Conversational Dialogue. AI Magazine 22, 4 (Dec. 2001), 39. https://doi.org/10.1609/aimag.v22i4.1591
- Koedinger et al. (2004) Kenneth R. Koedinger, Vincent Aleven, Neil Heffernan, Bruce McLaren, and Matthew Hockenberry. 2004. Opening the Door to Non-programmers: Authoring Intelligent Tutor Behavior by Demonstration. In Intelligent Tutoring Systems, James C. Lester, Rosa Maria Vicari, and Fábio Paraguaçu (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 162–174.
- Kojima et al. (2023) Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2023. Large Language Models are Zero-Shot Reasoners. arXiv:2205.11916 [cs.CL]
- Liu et al. (2023) Yi Liu, Gelei Deng, Zhengzi Xu, Yuekang Li, Yaowen Zheng, Ying Zhang, Lida Zhao, Tianwei Zhang, and Yang Liu. 2023. Jailbreaking ChatGPT via Prompt Engineering: An Empirical Study. arXiv:2305.13860 [cs.SE]
- Macina et al. (2023) Jakub Macina, Nico Daheim, Sankalan Pal Chowdhury, Tanmay Sinha, Manu Kapur, Iryna Gurevych, and Mrinmaya Sachan. 2023. MathDial: A Dialogue Tutoring Dataset with Rich Pedagogical Properties Grounded in Math Reasoning Problems. arXiv:2305.14536 [cs.CL]
- Nye et al. (2014) Benjamin Nye, Arthur Graesser, and Xiangen Hu. 2014. AutoTutor and Family: A Review of 17 Years of Natural Language Tutoring. International Journal of Artificial Intelligence in Education 24 (12 2014). https://doi.org/10.1007/s40593-014-0029-5
- Nye et al. (2021) Maxwell Nye, Anders Andreassen, Guy Gur-Ari, Henryk Witold Michalewski, Jacob Austin, David Bieber, David Martin Dohan, Aitor Lewkowycz, Maarten Paul Bosma, David Luan, Charles Sutton, and Augustus Odena. 2021. Show Your Work: Scratchpads for Intermediate Computation with Language Models. https://arxiv.org/abs/2112.00114
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL]
- Paladines and Ramirez (2020) José Paladines and Jaime Ramirez. 2020. A Systematic Literature Review of Intelligent Tutoring Systems With Dialogue in Natural Language. IEEE Access 8 (2020), 164246–164267. https://doi.org/10.1109/ACCESS.2020.3021383
- Phung et al. (2023) Tung Phung, Victor-Alexandru Pădurean, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023. Generative AI for Programming Education: Benchmarking ChatGPT, GPT-4, and Human Tutors. In Proceedings of the 2023 ACM Conference on International Computing Education Research - Volume 2 (Chicago, IL, USA) (ICER ’23). Association for Computing Machinery, New York, NY, USA, 41–42. https://doi.org/10.1145/3568812.3603476
- Rajala et al. (2023) Jaakko Rajala, Jenni Hukkanen, Maria Hartikainen, and Pia Niemelä. 2023. ”Call me Kiran” – ChatGPT as a Tutoring Chatbot in a Computer Science Course. In Proceedings of the 26th International Academic Mindtrek Conference. 83–94. https://dl.acm.org/doi/pdf/10.1145/3616961.3616974
- Reiser (2004) Brian Reiser. 2004. Scaffolding Complex Learning: The Mechanisms of Structuring and Problematizing Student Work. Journal Of The Learning Sciences 13 (01 2004), 273–304. https://doi.org/10.1207/s15327809jls1303_2
- Rich et al. (2018) Kathryn M. Rich, T. Andrew Binkowski, Carla Strickland, and Diana Franklin. 2018. Decomposition: A K-8 Computational Thinking Learning Trajectory. In Proceedings of the 2018 ACM Conference on International Computing Education Research (Espoo, Finland) (ICER ’18). Association for Computing Machinery, New York, NY, USA, 124–132. https://doi.org/10.1145/3230977.3230979
- Schmucker et al. (2023) Robin Schmucker, Meng Xia, Amos Azaria, and Tom Mitchell. 2023. Ruffle&Riley: Towards the Automated Induction of Conversational Tutoring Systems. arXiv:2310.01420 [cs.CL]
- Sleeman and Brown (1982) D. Sleeman and J.S. Brown. 1982. Intelligent Tutoring Systems. Academic Press. https://books.google.ch/books?id=pjqcAAAAMAAJ
- Song and Becker (2014) Ting Song and Kurt Becker. 2014. Expert vs. novice: Problem decomposition/recomposition in engineering design. In 2014 International Conference on interactive collaborative learning (ICL). IEEE, 181–190.
- Sonkar et al. (2023) Shashank Sonkar, Naiming Liu, Debshila Basu Mallick, and Richard G. Baraniuk. 2023. CLASS: A Design Framework for building Intelligent Tutoring Systems based on Learning Science principles. arXiv:2305.13272 [cs.CL]
- Yarkin and Coon (2022) Ferhat Yarkin and Justin P. Coon. 2022. Simple Gray Coding and LLR Calculation for MDS Modulation Systems. CoRR abs/2201.08237 (2022). arXiv:2201.08237 https://arxiv.org/abs/2201.08237
- Yu et al. (2023) Longhui Yu, Weisen Jiang, Han Shi, Jincheng Yu, Zhengying Liu, Yu Zhang, James T. Kwok, Zhenguo Li, Adrian Weller, and Weiyang Liu. 2023. MetaMath: Bootstrap Your Own Mathematical Questions for Large Language Models. arXiv:2309.12284 [cs.CL]
- Zhang et al. (2023) Muru Zhang, Ofir Press, William Merrill, Alisa Liu, and Noah A. Smith. 2023. How Language Model Hallucinations Can Snowball. arXiv:2305.13534 [cs.CL]
- Zhang et al. (2020) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q. Weinberger, and Yoav Artzi. 2020. BERTScore: Evaluating Text Generation with BERT. arXiv:1904.09675 [cs.CL]
| Hint | Recent history | Question | Solution Step | Rating |
Tell me this instead: What is the base number of books before Wendy adds the 20 extra books if she doubles the number of slots she has, which is 36? |
Tutor: You are correct upto where you say “Calculate the total number of slots in the shelving system by multiplying the number of rows by the number of columns. 6 rows * 6 columns = >36 slots.” What do you think should be the next step?
Wendy wants to place 20 more than double the number of books in a shelving system with 6 rows and 6 columns. How many books will she need to carry to complete her task? |
Double the number of slots to find the base number of books before adding the 20 extra books Wendy wants to place. 36 slots * 2 = <<36*2=72>>72 books. |
Grammatical:
YNN Relevant: YYY Not redundant: YYN Not revealing: YYY |
Tell me this instead: What is the profit when the total cost is subtracted from the new value of the house if the new value is $200,000 and the total cost is $130,000? |
Tutor: 200000 is correct, well done. What is the next step?
Josh decides to try flipping a house. He buys a house for $80,000 and then puts in $50,000 in repairs. This increased the value of the house by 150%. How much profit did he make? |
Find out the profit by subtracting the total cost from the new value of the house. $200,000-$130,000 = <<200000-130000=70000>>$70,000. |
Grammatical:
YYY Relevant: NNN Not redundant: NYY Not revealing: YNY |
Tell me this instead: What is the total amount of money earned in the second year from the sales of books if each was sold at the same price? |
Tutor: 65 is correct, well done. What is the next step?
Elise has been selling her Dad’s collection of 250 books for three years. Each book sells at 20$, and she sold twice as many books in the first year as she has sold in the current year. There are currently 50 unsold books, and her sales number this year is 45. What’s the total amount of money she earned in the second year? |
Compute the total amount of money earned in the second year by multiplying the number of books sold by the price per book: <<65 * 20 = 1300>>1300 dollars. |
Grammatical:
YYY Relevant: YYY Not redundant: NNN Not revealing: YYN |
Tell me this instead: What is the total cost per person for engaging in minigolf, purchasing arcade tokens, and enjoying two go-kart rides? |
Tutor:What should be the first step?
Morgan’s dad said that she had $90 budgeted for her birthday party. She wants to make sure she and her friends all get to play one round of mini-golf, have $5 in arcade tokens, and get to ride the go-karts twice. A round of mini-golf is $5. The Go-karts cost $10 a ride. How many friends can she invite? |
Calculate the total cost per person for mini-golf, arcade tokens, and two go-kart rides. 5 + 5 + (10 x 2) = <<5 + 5 + (10 * 2)=30>>30 dollars per person. |
Grammatical:
YYY Relevant: YYY Not redundant: YYY Not revealing: NYN |
| Hint | Recent history | Question | Solution Step | Rating |
That is not quite right. Here is a Hint: To calculate the total monthly expense for Madeline’s dog, you need to sum the individual costs for food, treats, and medicine. Consider each expense as a separate amount and add them together to find the total. Remember, the food expense is a two-digit number ending in zero, and both the treat and medicine expenses have the same amount, which is also a two-digit number ending in zero |
Tutor: Tell me this instead: What is the total monthly expense for Madeline’s dog when combining the costs of food, treats, and medicine?
Every month, Madeline has to buy food, treats, and medicine for her dog. Food costs $25 per week. Treats cost $20 per month. Medicine costs $100 per month. How much money does Madeline spend on her dog per year if there are 4 weeks in a month? |
Add the monthly costs of food, treats, and medicine to find the total monthly expenses for Madeline’s dog. $100 (food per month)+ $20 (treats per month) +$100 (medicine per month) = 100+20+100=220$220 per month |
Grammatical:
YNY Specific: YNN Helpful: YNY Not revealing: YYY |
That is not quite right. Here is a Hint: To determine the number of games Bobby has, first identify Brian’s games and multiply that amount by three. Once you have that number, decrease it by five and you will find out Bobby’s total games. |
Tutor: Tell me this instead: What is the number of games Bobby has if you take away 5 from three times the number of Brian’s games?
Brian’s friend Bobby has 5 fewer than 3 times as many video games as Brian does. If Brian has 20 video games but lost 5 right before the comparison was made, how many does Bobby have? |
Subtract 5 from the tripled number of Brian’s games to find out how many games Bobby has.
Grammatical:
YYY Specific: YYY Helpful: YYY Not revealing: NNY |
An example solution and utterance space from the MultiArith dataset obtained using GPT4

### v2

@@
 Large Language Models (LLMs) have found several use cases in education, ranging from automatic question generation to essay evaluation.
-In this paper, we explore the potential of using Large Language Models (LLMs) to author Intelligent Tutoring Systems.
-A common pitfall of LLMs is their straying from desired pedagogical strategies such as leaking the answer to the student, and in general, providing no guarantees.
-We posit that while LLMs with certain guardrails can take the place of subject experts, the overall pedagogical design still needs to be handcrafted for the best learning results.
+In this paper, we explore the potential of using LLMs to author Intelligent Tutoring Systems.
+A common pitfall of using LLMs as tutors is their straying from desired pedagogical strategies such as leaking the answer to the student, and in general, providing no guarantees on the validity or appropriateness of the tutor assistance.
+We argue that while LLMs with certain guardrails can take the place of subject experts, the overall pedagogical design still needs to be handcrafted for the best learning results.
 Based on this principle, we create a sample end-to-end tutoring system named MWPTutor, which uses LLMs to fill in the state space of a pre-defined finite state transducer.
 This approach retains the structure and the pedagogy of traditional tutoring systems that has been developed over the years by learning scientists but brings in additional flexibility of LLM-based approaches.
-Through a human evaluation study on two datasets based on math word problems, we show that our hybrid approach achieves a better overall tutoring score than an instructed, but otherwise free-form, GPT-4.
-MWPTutor is completely modular and opens up the scope for the community to improve its performance by improving individual modules or using different teaching strategies that it can follow.
+Through a human evaluation study on two datasets with math word problems, we show that our hybrid approach achieves a better overall tutoring score than an instructed, but otherwise free-form, GPT-4.
+MWPTutor is completely modular and opens up the scope for the community to improve its performance by refining its individual modules or using different teaching strategies that it can follow.
 MWPTutor
@@
 2.2. Large Language Models
-Large Language Models (LLMs) are Transformer-based models that serve as general-purpose tools in natural language processing (Bommasani et al., 2022).
+Large Language Models (LLMs) are Transformer-based models that serve as general-purpose tools in natural language processing (Bommasani et al., 2021).
 With recent improvements in LLMs and the availability of LLM-as-a-service models like GPT (OpenAI, 2023) and BARD (Yarkin and Coon, 2022), it has become easy to use LLMs as tutors by prompting them with the desired qualities of a tutor and then letting them interact with a student.
-Some works in the programming domain like GPTutor (Chen et al., 2023) and TestGPT (Phung et al., 2023) attempt to use GPT4 similar to information delivery systems.
+Some works in the programming domain like GPTutor (Chen et al., 2023) and Phung et al.
+(Phung et al., 2023b, a) attempt to use GPT4 similar to information delivery systems.
 There have also been systems designed to function as Dialog agents, like Ruffle&Riley (Schmucker et al., 2023) which includes multiple GPT4 agents functioning with different prompts, and SPOCK (Sonkar et al., 2023) which attempts to distill out tutoring knowledge from GPT4 into a smaller LLM.
@@
 4.4. Answer and Path Detection
-To execute the pedagogical strategies described above, we need to be able to detect if the student has given a correct answer.
+To execute the pedagogical strategies described above, we need to detect if the student has given a correct answer.
 Further, the hint state should only trigger if the student fails to propose a correct follow-up.
@@
 MultiArith consists of simple Math Word Problems that are usually solvable in 1 to 3 steps.
-GPT3 was already able to solve over 90% of these problems with the right prompts (Kojima et al., 2023), and in our tests, we found that GPT4 was able to get a perfect score in the MWP solver setting.
+GPT3 was already able to solve over 90% of these problems with the right prompts (Kojima et al., 2022), and in our tests, we found that GPT4 was able to get a perfect score in the MWP solver setting.
 With this dataset, we want to test if we can achieve the same performance in the interactive tutoring setting.
@@
 These issues can be fixed by better LLMs and more sophisticated techniques outlined in future work.
-Several potential steps have already been outlined in Section 4, and we hope to soon have better versions of MWPTutor and other similar systems that can realize one-on-one tutoring at scale.
+Several potential steps have already been outlined, and we hope to soon have better versions of MWPTutor and other similar systems that can realize one-on-one tutoring at scale.
 Acknowledgements.
@@
 Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S.
-Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, Shyamal Buch, Dallas Card, Rodrigo Castellon, Niladri Chatterji, Annie Chen, Kathleen Creel, Jared Quincy Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren Gillespie, Karan Goel, Noah Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E.
-Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, Omar Khattab, Pang Wei Koh, Mark Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D.
-Manning, Suvir Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Ben Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, Julian Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Rob Reich, Hongyu Ren, Frieda Rong, Yusuf Roohani, Camilo Ruiz, Jack Ryan, Christopher Ré, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishnan Srinivasan, Alex Tamkin, Rohan Taori, Armin W.
+Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, S.
+Buch, Dallas Card, Rodrigo Castellon, Niladri S.
+Chatterji, Annie S.
+Chen, Kathleen A.
+Creel, Jared Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren E.
+Gillespie, Karan Goel, Noah D.
+Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E.
+Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas F.
+Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, O.
+Khattab, Pang Wei Koh, Mark S.
+Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D.
+Manning, Suvir P.
+Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Benjamin Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, J.
+F.
+Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Robert Reich, Hongyu Ren, Frieda Rong, Yusuf H.
+Roohani, Camilo Ruiz, Jack Ryan, Christopher R’e, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishna Parasuram Srinivasan, Alex Tamkin, Rohan Taori, Armin W.
 Thomas, Florian Tramèr, Rose E.
-Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang.
-2022.
+Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei A.
+Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang.
+2021.
 On the Opportunities and Risks of Foundation Models.
-arXiv:2108.07258 [cs.LG]
+ArXiv (2021).
+https://crfm.stanford.edu/assets/report.pdf
 - Cai et al.
@@
 Springer Berlin Heidelberg, Berlin, Heidelberg, 162–174.
-- Kojima et al. (2023) Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2023. Large Language Models are Zero-Shot Reasoners. arXiv:2205.11916 [cs.CL]
+- Kojima et al.
+(2022) Takeshi Kojima, Shixiang (Shane) Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa.
+2022.
+Large Language Models are Zero-Shot Reasoners.
+In Advances in Neural Information Processing Systems, S.
+Koyejo, S.
+Mohamed, A.
+Agarwal, D.
+Belgrave, K.
+Cho, and A.
+Oh (Eds.), Vol.
+35.
+Curran Associates, Inc., 22199–22213.
+https://proceedings.neurips.cc/paper_files/paper/2022/file/8bb0d291acd4acf06ef112099c16f326-Paper-Conference.pdf
 - Liu et al.
@@
 MathDial: A Dialogue Tutoring Dataset with Rich Pedagogical Properties Grounded in Math Reasoning Problems.
-arXiv:2305.14536 [cs.CL]
+In The 2023 Conference on Empirical Methods in Natural Language Processing.
+https://openreview.net/forum?id=fyza2OQ9NI
 - Nye et al.
@@
 - Phung et al.
-(2023) Tung Phung, Victor-Alexandru Pădurean, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares.
-2023.
+(2023a) Tung Phung, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares.
+2023a.
+Generating High-Precision Feedback for Programming Syntax Errors using Large Language Models.
+In Proceedings of the 16th International Conference on Educational Data Mining, Mingyu Feng, Tanja Käser, and Partha Talukdar (Eds.).
+International Educational Data Mining Society, Bengaluru, India, 370–377.
+https://doi.org/10.5281/zenodo.8115653
+- Phung et al.
+(2023b) Tung Phung, Victor-Alexandru Pădurean, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares.
+2023b.
 Generative AI for Programming Education: Benchmarking ChatGPT, GPT-4, and Human Tutors.

### v3

@@
-Scaling the Authoring of AutoTutors with Large Language Models
+Scaling the Authoring of AutoTutors with Large Language Models Large Language Models with Rich Pedagogy and GuardRails AutoTutor meets Large Language Models:
 Abstract.
@@
 This is typically done only once for an entire domain.
-The second stage, performed by either teachers or domain experts, is to author the scripts111see Appendix A in (Graesser et al., 2000) for an example script, which needs to be done separately for every topic, question, or concept.
+The second stage, performed by either teachers or domain experts, is to author the scripts111See Appendix A in (Graesser et al., 2000) for an example script., which needs to be done separately for every topic, question, or concept.
 In the standard framework, both these stages are done by humans.
 In contrast, in the LLM-only framework discussed in the previous paragraph, both these stages are carried out (implicitly or explicitly) by LLMs.
-In this paper, we consider a mixed approach where we only delegate the job of domain experts to an LLM, while the first stage is still left to learning experts.
+In this paper, we develop a mixed approach where we only delegate the job of domain experts to an LLM, while the first stage is still left to learning experts.
 This idea has previously been hinted at by Cai et al.
 (Cai et al., 2019), who propose using a question-answering system to automatically list potential misconceptions for a given topic.
-We extend this idea to have an LLM author the entire script for the tutoring system, which is accurate enough to be deployed without human validation but is fully open to human validation.
-Our contributions in this paper are as follows:
-We adapt the Dialog Based tutoring framework used by ITS’s to make it possible for a Large Language Model to author it.
-We design MWPTutor, a tutor for math word problems based on the above principles.
-We compare the performance of MWPTutor with that of GPT4 being used as a tutor both quantitatively and qualitatively.
+We extend this idea to have an LLM author the entire script for the tutoring system.
+This authoring process includes internal checks and guardrails, making the results accurate enough to be deployed without human validation.
+However, should validation be essential, we also describe a version that can be checked by humans to increase trust.
+In our experiments, we find that our system is better at tutoring an AI-simulated student than instructed, but otherwise free-form GPT4.
+We also find that human evaluators judge that individual utterances are indeed following the specified pedagogy.
+However, manual analysis of the conversations shows that there is still some scope for improvement.
+In summary, our contributions in this paper are as follows:
+We adapt the Dialog Based tutoring framework used by ITS’s to make it possible for an LLM to author it.
+We design MWPTutor, a tutor for math word problems which can function both with or without live calls to an LLM based on the above principles.
+We compare the performance of both versions of MWPTutor with that of GPT4 being used as a tutor both quantitatively and qualitatively.
 2. Background
@@
 The idea was to roleplay a human tutor with a computer program that could interact with the student and fulfill the requirement of personalized tutoring.
-The computer program may not fully replace human tutors but often provides an affordable and scalable addition.
+This may not fully replace human tutors but often provides an affordable and scalable addition.
 An important aspect that differentiates ITSs from simply showing the students the problem solution is their ability to provide step-by-step feedback which promotes long-term learning as opposed to short-term problem-solving success.
 This was often implemented using rules that check for certain errors using string-matching and act on them with pre-defined templates.
-For example, LISP Tutor (Anderson et al., 1985) contains rules to account for common programming errors made by students.
+For example, LISP Tutor (Anderson et al., 1985) uses rules to account for common programming errors made by students.
 With advancements in AI, ITSs transitioned to dialog as a method of tutoring, moving ITSs from being information-delivery systems to dialog-based agents.
@@
 Not knowing the student misconceptions in advance, domain experts have to come up with rules for numerous possibilities, introducing a different set of scaling issues.
-One hour of interaction with a student costs over 100 hours of domain expert time to author the ITS (Koedinger et al., 2004).
+One hour of interaction with a student costs over hours of domain expert time to author the ITS (Koedinger et al., 2004).
 This is in addition to the time spent by learning scientists designing the pedagogy and dialog strategies.
@@
 The optimal tutoring support differs between the two examples.
-While the math word problem starts with subquestions to compute intermediate quantities, answering the history questions first requires providing historical context.
 3.1. Solution Decomposition
@@
 3.3. Scaffolding the LLM
-While modern LLMs fine-tuned with RLHF are quite good at following instructions, they can at times struggle with more complex tasks.
+While modern LLMs fine-tuned to follow natural language instructions are quite good at doing so, they can at times struggle with more complex tasks.
 Because of the probabilistic nature of LLMs, reattempting the same problem can lead to different answers.
 To take the best advantage of these facts, we want to have automated checks for every LLM turn in the strategy space, so calls can be repeated until some checks have been met.
-For example, D3 could require that the numerical answer (450 for A1) is included in the utterance, while P2 could require the missing keywords to be absent.
+For example, D3 could require that the numerical answer ( for A1) is included in the utterance, while P2 could require the missing keywords to be absent.
 These checks can take different forms depending on the domain, like limiting the length of an utterance or avoiding the use of more complex language.
-3.4. Control and Reliability
-One of the primary goals of having LLMs author a state space rather than having them interact directly with the students is that in this way, one can examine and trust the model without having to necessarily trust the LLM used to generate it.
-Based on this, we store and make it possible to inspect beforehand the entire utterance space.
-Changes to the utterances themselves can be made on-the-go, e.g.
-if a teacher feels it is better to split our example in P2 into two sentences, they can simply edit it to say “Who was assassinated?” and “Who carried out the assassination” with no further changes required.
-Changes to the solution space, however, need the utterances to be regenerated.
-Caching the utterance LLMs like GPT also makes the system immune to changes made to GPT in terms of availability, cost, or rate limits.
-For open LLMs, this allows for the use of Larger LLMs which not every user can run locally.
-The most important benefit, however, is the potential for teachers/domain experts to edit the state space should they want, therefore allowing us to combine human expertise with machine efficiency.
-Live calls to an LLM can often help and can be incorporated into the state space.
-For example, in the MWP we could make a call to GPT to take over a step in between D2 and D3 if we want to minimize the chance of D3.
-Should live calls be implemented, however, there should ideally be non-LLM fallbacks ie we should have cached and/or procedural alternatives available should a call to an LLM not be possible at deployment time.
+3.4. The Trade-off between Reliability and Contextual Awareness
+Once we have prompts for each utterance and relevant checks, we can query an LLM to generate the tutor utterances.
+Here we have a choice on when the generation is done.
+One option is to generate all possible utterances at design time, and storing them into the state space.
+This would be similar to what traditional ITSs would do.
+Doing this not only allows the tutor to run without an active internet connection, but having an exhaustive set of utterances stored allows stakeholders to trust the tutor without necessarily trusting the LLM used to generate the utterances.
+This is possible because one can reliably say that there would be no “new” errors or hallucinations since the LLM is not used after the design stage.
+Despite these advantages, pre-generating means that all possible scenarios that can arise during the conversation need to be anticipated in advance, which is extremely hard to achieve.
+Therefore, such a model is likely to feel somewhat oblivious of the context of the conversation, which might reduce student engagement, which in turn, may reduce learning gains (Martin and Bolliger, 2018; Bryson and Hand, 2007; Parsons and Taylor, 2011) The alternative to this, is to generate the utterances during the conversation.
+This allows us to include recent utterances from both the tutor and the student into the prompt, thereby generating a more context-aware response.
+Due to changing context, it is also possible to allow multiple utterances for a single state, allowing for more fine-grained feedback.
+Doing this, however, leaves the LLM open to hallucinations and jailbreaks as discussed earlier.
+The choice we make depends of factors like how good we believe the backend LLMs are, how much the policymakers trust them, how complex is the problem etc.
+We do not believe that any of these options is better than the other, and the best option might be a combination of both.
 4. MWPTutor
@@
 However, to not become constrained to the availability of such solutions, and to allow diverse solution paths, we restrict MWPTutor to only the final numerical answer to the problem, and generate the solutions ourselves.
+Further, to highlight the tradeoff mentioned in section 3.4, we shall design two slightly different versions MWPTutor, which is allowed to make live calls to the LLM, and MWPTutor which is not.
+Note that unless otherwise specified, designs mentioned in this section apply to both versions.
 symbol indicates student utterance inputs, while the
@@
 These are preceded by a “pump”, which is the generic utterance “What should be the next step?” The overall strategy then proceeds as follows:
-| Pump |
-asks the student to propose the next step. If the student proposes the step correctly, we ask them to solve it. |
-| Hint |
-is only triggered if the student fails to propose a valid next step, or fails to execute the step they propose.
+| Pump | asks the student to propose the next step. If the student proposes the step correctly, we ask them to solve it. |
+| Hint | is only triggered if the student fails to propose a valid next step, or fails to execute the step they propose.
 It essentially gives out the step and asks the student for the calculation.
@@
 |
-| Prompt |
-is triggered if the student fails to execute the step after the hint. This should push the student closer to the answer, without actually revealing the answer. |
-| Assertion |
-is triggered if the student fails to answer even after the prompt. This should simply give out the answer, and have the student acknowledge it. |
-In our running example, hints, prompts, and assertions would correspond to D1, D2, and D3 respectively. The pump is not included.
+| Prompt | is triggered if the student fails to execute the step after the hint. This should push the student closer to the answer, without actually revealing the answer. |
+| Assertion | is triggered if the student fails to answer even after the prompt. This should simply give out the answer, and have the student acknowledge it. |
+In our running example, hints, prompts, and assertions correspond to D1, D2, and D3 respectively. The pump is not included.
 For scaffolding the LLM, both hints and prompts are required to avoid explicitly outputting the right-hand side of the equation in the step, with the hint additionally having to avoid saying any words that indicate arithmetic operators, currently “add, subtract, sum, multiply, product, divide, quotient.” The assertion, on the other hand, must contain the right-hand side.
 For pumps, there is no use of an LLM, so we do not need a check.
+For MWPTutor, we allow up to prompts before moving to assertion.
+For both hints and prompts, the LLM is given the steps the student has correctly solved so far and the last utterances in the dialogue ( each for the student and the tutor) in addition to the solution step and strategy (see Section A.3 for details)
 4.4. Answer and Path Detection
@@
 For this, we need to have a way to match paths.
-We run into the same matching problems described in Section 4.2, and resort to similar basic techniques.
+We run into the same matching problems described in Section 4.2, and in case of MWPTutor, resort to similar basic techniques.
 The correct answer detection algorithm is shown in Algorithm 2.
@@
 If multiple steps fulfill this, or no step does, we pick the one with the best BertScore alignment to the student utterance.
-The fact that we currently ignore all non-numeric content in the student response is one of the major limitations in our work, as it can miss a large number of “correct” utterances from the student, but we defer the resolution of this to future work.
+Since MWPTutor is allowed to make live calls to the LLM, we make use of it for both correct answer detection, and picking the best path to take when in the “pump” stage.
+For correct answer detection, we use the RHS match as an additional criterion, while for picking the best path, we rely entirely on the LLM’s judgement.
+The fact that MWPTutor currently ignores all non-numeric content in the student response is one of its major limitations, as it can miss a large number of “correct” utterances from the student, but we defer the resolution of this to future work.
 Inputs: CurrentSteps from Solution step space, and Student Utterance
@@
 To obtain to full finite state transducer structure, one must roll out the states along the solution space.
-Excluding the unlikely case where the student fails to repeat the answer after the assertion state, the overall conversation should not last more than utterance pairs, where is the number of steps.
+Excluding the unlikely case where the student fails to repeat the answer after the assertion state, the overall conversation should not last more than utterance pairs for MWPTutor, and for MWPTutor, where is the number of steps.
 5. Experiments
@@
 5.1.2. Models.
-At the time of testing, GPT4555gpt-4-1106-preview queried in January 2024 was considered to be one of the best-performing LLMs, so we use it both as a backend for MWPTutor and as the baseline single prompt LLM.
+At the time of testing, GPT4555gpt-4-1106-preview queried in January 2024 was considered to be one of the best-performing LLMs, so we use it both as a backend for the two MWPTutor’s and as the baseline single prompt LLM.
 To best align the single prompt GPT4 with the MWPTutor doctrine, we ask it to roleplay a tutor, setting the primary objective of getting to the correct answer as the goal of the conversation, and the secondary objective of not revealing the answer as a personality trait.
 Further, to test the potential of an open-source LLM as a backend, we have a second version of MWPTutor with a MetaMath-13B (Yu et al., 2023).
-MetaMath is a version of Meta AI’s open source LLM LLaMa666https://llama.meta.com/ fine-tuned to solve math word problems, and is, to the best of our knowledge, the best open LLM for solving math word problems.
-We refer to these three models as MWPTutor, GPT4, and MWPTutor respectively.
+MetaMath is a version of Meta AI’s open source LLM LLaMa666llama.meta.com fine-tuned to solve math word problems, and is, to the best of our knowledge, the best open LLM for solving math word problems.
+However, since MWPTutor would require the Backend to run throughout the conversation, which is unsustainable for a locally-run model, we only use this with MWPTutor.
+We shall refer to the live model as MWPTutor, the cached models as MWPTutor and MWPTutor based on the backend, and the single prompt model simply as GPT4 hereafter.
 5.1.3. Datasets.
@@
 All models are able to transfer their near-perfect solving scores to near-perfect tutoring scores.
-GPT4 performs slightly worse, although it has a telling rate.
-Both versions of MWPTutor have a perfect success rate.
+Both GPT4 and MWPTutor have no telling, while the telling rate of the MWPTutor models is also almost negligble.
+All versions of MWPTutor have a perfect success rate, with GPT4 not far behind.
+These results demonstrate that all our models are aligned with our doctrine, and are able to follow it well on a simpler dataset.
 It must be noted that the metric is noisy since it only requires the number to match, with no restriction on the context, e.g.
@@
 Nevertheless, such errors are unlikely to be frequent, and all models work reasonably well on these problems.
-| MWPTutor | GPT4 | |||
-| Metric | LLaMa | GPT | ||
-MultiArith |
-Success Rate | 1.00 | 1.00 | 0.98 |
-| Telling Rate | 0.01 | 0.01 | 0.00 | |
-| Adjusted Success | 0.99 | 0.99 | 0.98 | |
-| Tutoring Score | 0.99 | 0.99 | 0.98 | |
-MathDial |
-Success Rate | 0.91 | 0.96 | 0.82 |
-| Telling Rate | 0.33 | 0.11 | 0.00 | |
-| Adjusted Success | 0.58 | 0.84 | 0.82 | |
-| Tutoring Score | 0.71 | 0.90 | 0.82 | |
+| MWPTutor | MWPT. | GPT4 | |||
+| LLaMa | GPT4 | GPT4 | |||
+| MultiArith | Success Rate | 1.00 | 1.00 | 1.00 | 0.98 |
+| Telling Rate | 0.01 | 0.01 | 0.00 | 0.00 | |
+| Adjusted Success | 0.99 | 0.99 | 1.00 | 0.98 | |
+| Tutoring Score | 0.99 | 0.99 | 1.00 | 0.98 | |
+| MathDial | Success Rate | 0.91 | 0.96 | 1.00 | 0.82 |
+| Telling Rate | 0.33 | 0.11 | 0.00 | 0.00 | |
+| Adjusted Success | 0.58 | 0.84 | 1.00 | 0.82 | |
+| Tutoring Score | 0.71 | 0.90 | 1.00 | 0.82 | |
 For the harder MathDial dataset, we wanted to avoid matching errors as described earlier, and so have human annotators mark out utterances where the student or the tutor says the final answer.
 Each conversation was annotated by annotators, and we took the majority in cases where they disagreed.
-Table 1 summarises our result.
-While GPT4 maintains its 0% telling rate, MWPTutor gets more problems correct overall, and more when adjusted for telling.
+The bottom half of Table 1 summarises our result.
+Both GPT4 and MWPTutor maintain their 0% telling rate, but MWPTutor is the only model able to maintain its 100% success rate.
+This clearly establishes MWPTutor as the model of choice should live access to GPT4 not be an issue, as discussed in Section 3.4.
+Of the cached models, MWPTutor gets more problems correct than GPT4, and more when adjusted for telling.
 This gives MWPTutor an advantage on GPT4 in terms of the final score.
 MWPTutor also gets a higher success rate than GPT4, which is somewhat surprising since its backend model is significantly weaker than GPT4 in the MWP Solver setting (Yu et al., 2023).
-This increased success rate however comes at a cost of a very high telling rate, giving it a significantly poorer tutoring score overall, so we do not recommend the use of MetaMath as the backend model unless GPT4-like models need to be avoided for some external reason.
+This increased success rate however comes at a cost of a very high telling rate, giving it a significantly poorer tutoring score overall.
+Based on this we do not recommend the use of MetaMath as the backend model unless GPT4-like models need to be avoided for some external reason.
 One thing worth mentioning here is that once problematic questions have been identified for either version of MWPTutor, one can have human annotators redo only those questions, thereby increasing the success rate, something that is not possible for GPT4.
+MWPTutor is somewhere in the middle, as we can only “fix” the solution step space.
 5.3. Utterance Evaluation
 A major shortcoming of AI-on-AI-tutoring is that the AI student may not be affected by the sometimes subpar tutor responses.
-This is more common in MWPTutor than in GPT4, as the former can only pick from a pre-defined set of responses and not create new utterances on the fly in response to finer details in the student requirements.
-To understand if this is a major issue, we hired human evaluators who worked in the education sector and asked them to evaluate utterances made by MWPTutor for MathDial.
+To better understand the quality of utterances both in terms of them executing the pedagogical strategy and in term of their context awareness, we hired human evaluators who worked in the education sector and asked them to evaluate utterances made by MWPTutor and MWPTutor for MathDial.
 Because the pumps are generic and very similar to each other and assertions essentially boil down to reading out the answers, we focus only on the hints and prompts in our evaluation.
-Across the conversations, MWPTutor and the student model went over 135 solution steps.
-Of these went to the hint state, while reached the prompt state.
+Across the conversations, MWPTutor and the student model went over solution steps Of these went to the hint state, while reached the prompt state.
+In the case of MWPTutor, we had steps total, of which went to a prompt hint state, went to the prompt stage.
+of these would then use a second prompt and would use a third, giving us a total of prompts.
 We presented each of these utterances to human evaluators, with the past conversation, the original math word problem, and the current solution step as context.
 We boil down the requirements set on hints and prompts set in Section 4.3 to the following binary criteria which we asked the evaluators to evaluate the utterances on:
-Grammatical: The utterance is linguistically and grammatically correct.
-Not Reveal: The utterance does not reveal the final answer for the step.
-Relevant (hints only): The hint clearly addresses the step, ie, the only way to correctly answer the question is to complete the current step.
-Not Redundant (hints only): The hint is not redundant to what the student already said in the previous step ie it makes sense to ask this question after the previous response.
-Specific (prompts only): The prompt is specific to the step, and not just a vague lead-on statement.
-Helpful (prompts only): The prompt can potentially help the student make progress in the current step, given the previous student’s utterance.
+Grammatical: Is linguistically and grammatically correct.
+Not Reveal: Does not reveal the final answer for the step.
+Relevant (hints only): Addresses the step and the only way to correctly answer the question is to complete the current step.
+Not Redundant (hints only): Is not redundant to what the student already said in the previous step ie it makes sense to ask this question after the previous response.
+Specific (prompts only): Is specific to the step, and not just a vague lead-on statement.
+Helpful (prompts only): Can help the student make progress in the current step, given the previous student’s utterance.
 The annotators were paid the Prolific suggested wage of 10 GBP/hour.
-We show the annotation user interface in Figure 6 and some example annotations in Table 2 and Table 3.
+We show the annotation user interface in Figure 6 and some example annotations in Tables 2, LABEL:, 3, LABEL:, 4, LABEL: and 5.
 Figures 4 and 5 summarise the responses from the evaluators.
-Almost all utterances are considered Grammatical by the evaluators, which is not surprising given they were generated by GPT4.
-Most of the utterances are also marked to be non-revealing, which is also not surprising given that any proposed utterances that directly revealed the answer were assigned to be regenerated.
+Almost all utterances are considered Grammatical by the evaluators, which is in line with the known grammatical prowess of GPT4.
+It is unclear why there is a slightly higher count of non-grammatical utterances in MWPTutor, but it could be potentially due to annotators finding the more eloquent responses of GPT4 harder to judge grammatically.
+Most of the utterances are also marked to be non-revealing, which is expected given that any proposed utterances that directly revealed the answer were assigned to be regenerated.
+MWPTutor performs slightly better here.
 The evaluators further seem to agree that most hints were relevant to the step and that most prompts are specific, implying that they do what is appropriate given the current solution step.
+Again, MWPTutor has an edge in this aspect.
+Of the two of its prompts found to be non-specific, one is a second prompt and one is a first prompt, so all third prompts are considered specific
 The biggest potential weakness of MWPTutor comes in the form of its utterances making sense in the context of previous student utterances due to issues discussed in Section 4.4, which is where the final two metrics come in.
-In this regard, all prompts are considered to be helpful by a majority of evaluators which is a great sign.
-Hints, however, are found to be redundant in almost a third of cases, showing that handling of student’s responses to pumps could be better.
-The current heuristic of matching the left-hand side is too demanding, and highly likely to ignore the student even if they propose a valid next step to the problem, which can break student immersion.
-Note that of the steps did not proceed to the hint state, implying more than half of the correct proposals by the student are being identified as such.
-Further, in a vast majority of cases where a correct proposal is misidentified, there is disagreement between human evaluators indicating there might be some merit in the rejection, though this is also the case for a majority of correct rejections.
+In this regard, all prompts from MWPTutor are considered to be helpful by a majority of evaluators which is a great sign.
+Surprisingly, MWPTutor ends up giving 5 prompts that are not considered helpful.
+One reason why this happens could be the fact that there are more prompts allowed for each step, making individual prompts less helpful.
+Regardless, it is still a small number.
+Hints, however, are found to be redundant in almost a third of cases for MWPTutor, showing that handling of student’s responses to pumps could be better.
+MWPTutor brings this number much lower, but still with redundant Hints, it is the second largest red bar in the graph, only behind its cached counterpart.
+It is clear that MWPTutor’s heuristic of matching the left-hand side is too demanding, and highly likely to ignore the student even if they propose a valid next step to the problem, which can break student immersion.
+It should be noted, however that of the steps did not proceed to the hint state for MWPTutor, implying more than half of the correct proposals by the student are being identified as such.
+Curiously, the number of non-redundant hints is almost the same for the two models.
+Further, in a vast majority of cases where a correct proposal is misidentified, there is disagreement between human evaluators indicating there might be some merit in the rejection, though this is also the case for a many of correct rejections.
 Overall these evaluations show that utterances generated for MWPTutor by the GPT4 backend are consistent with the strategy definitions, and the transition function uses them reasonably well.
+MWPTutor further increases the efficacy of the transition function.
+Both models, however, have scope for improvement.
 6. Qualitative Assessment of Weaknesses
-While both GPT4 and MWPTutor do quite well on quantitative metrics as seen above, these metrics do not fully grasp what it takes to make a good tutor.
-To better understand the preparedness of our models, we went over all conversations that happened on the MathDial dataset, for both GPT4 and MWPTutor and summarize our findings here.
+While all our models do quite well on quantitative metrics as seen above, these metrics do not fully grasp what it takes to make a good tutor.
+To better understand the preparedness of our models, we went over all conversations that happened on the MathDial dataset, for GPT4 and both MWPTutor models, and summarize our findings here.
 Firstly, MWPTutor tends to sound monotonous and dry.
@@
 Affective statements, like encouragements, challenges, etc.
-would further require accurate judgment of student confidence and correctness.
-In both these cases, we made a design choice to avoid the additional module, as they would create major failure nodes.
-GPT4 on the other hand, sounds a lot more natural, including finer feedback and probing, and even affective statements, although they are exclusively positive.
-Further, since it matches correct answers with limited attention to context, MWPTutor can sometimes get false positives, and inadvertently reveal the answer. Consider the following excerpt:
+would further require accurate judgment of student confidence and correctness, which we chose to avoid due to it being a possible failure node.
+GPT4 and MWPTutor sound a lot more natural, including finer feedback and probing, and even affective statements, although they are exclusively positive.
+The ability of MWPTutor in achieving appropriate affect statements might indicate that we might have been overzealous in not having affective statements for MWPTutor, and a similar effect might be achievable using a smaller language model.
+Further, since it matches correct answers with limited attention to context, MWPTutor can sometimes get false positives, and inadvertently reveal the answer. Consider the following:
 Here, the student gives the wrong answer of , but MWPTutor mistakes the intermediate calculation of as the correct answer, and effectively reveals it to the student.
 GPT4 is able to understand context usually, although it is far from perfect.
-As we discuss later, GPT4 fails much worse on this question too.
-While GPT4 conversations flow a lot better and sound human-like, it has a different set of issues.
-For one, GPT4 tends to overcomplicate questions or make redundant calculations.
+GPT4 fails much worse on this question too.
+MWPTutor, on the other hand, tends to have a false negatives, perhaps because needing approval from both the LLM and the RHS checker is a bit too strict.
+Further, unlike GPT4, it still cannot account for student solving multiple steps in one utterances, so at times it might look like it is holding the student back.
+In general, it handles the false negatives well, by either asking the student to repeat or revise steps or asking for clarifications, but ideally, we do not want to prolong conversations anymore than needed.
+While it does not have the one-step per utterance restriction, GPT4 still tends to overcomplicate questions or make redundant calculations.
 In one question, where the student failed to convert seconds into minutes, GPT4 gets the student to convert the seconds to hours first, and then back to minutes.
@@
 The conversation in Figure 1 is also an example of where this happens in the simpler MultiArith dataset.
-Since it explores multiple paths and is also limited in the steps it can take, MWPTutor is less likely to run into this issue, and handles the first two cases mentioned here pretty well, though it makes the same error in the last case.
-The biggest problem with GPT4, however, continues to be that GPT4 can lead the student to a wrong answer, sometimes even leading them away from the right answer.
+Due to its more contained and well-structured state space, MWPTutor is less likely to run into this issue, and handles the first two cases mentioned here pretty well, though it makes the same error in the last case.
+The biggest problem with GPT4, continues to be that GPT4 can lead the student to a wrong answer, sometimes even leading them away from the right answer.
 In all cases where GPT4 fails to get the student to the correct answer, it outright accepts a wrong answer as correct.
-Of MWPTutor’s failures one was a result of inadvertently revealing the final answer, while the other seems to have been marked as incorrect by the annotators because the tutor resorts to revealing rather quickly.
 In a further conversation, including the one regarding chicken feed described earlier, GPT4 misdirects the student onto the wrong path but somehow manages to get them to say the correct answer.
@@
 It must be reiterated here that the final correct answer is given to GPT4 as part of its prompt.
-MWPTutor by design will never okay a wrong answer.
-The only place where it does go wrong, it accepts an answer of instead of , but this can be chalked down to the fact that the question was asking for the “difference” between two quantities, and did not make it clear whether the sign was required (the first quantity was smaller.
+All MWPTutor models by design will never okay a wrong answer.
+Of MWPTutor’s failures one was a result of inadvertently revealing the final answer, while the other seems to have been marked as incorrect by the annotators because the tutor resorts to revealing rather quickly.
+Beyond these, in one question it accepts an answer of instead of , but this can be chalked down to the fact that the question was asking for the “difference” between two quantities, and did not make it clear whether the sign was required (the first quantity was smaller.
 Also GPT4 makes the same error.
 Both are marked as correct by annotators).
-It must, however, be noted that since it also relies on GPT4 for solution generation, MWPTutor can hallucinate, even if it did not happen in any of these conversations to the best of the authors’ judgement (We attempted to have the evaluators annotate the existence of wrong methods, but found too much disagreement in their responses to consider them viable).
-If this happens, however, one can fix them by changing the tree the first time they get noticed, something that is not possible with GPT4.
+We would like to clarify that, all MWPTutor models, especially MWPTutor can also hallucinate, even if it did not happen in any of these conversations to the best of the authors’ judgement (We attempted to have the evaluators annotate the existence of wrong methods, but found too much disagreement in their responses to consider them viable).
 Overall, we conclude that despite good showing on metrics, both methods are still lacking in quality, and decide to not test them on real students until we can significantly improve them.
@@
 Adding a richer structure to the Solution graph could vastly reduce the size of the solution space.
-In addition to this, each step of the solution should be made to follow a more standardized format.
+In addition, each step of the solution should be made to follow a more standardized format.
 Having a more precise definition of the step would allow us to use LSA-like techniques to improve our correct answer detection.
@@
 7.2. Improving the Strategy Space
-The current strategy space is linear and does not allow for very targeted reactions to the student utterances or the original student solution.
+The current strategy space for the cached models is very linear, and cannot react well to exact student responses.
 Once we have a better solution step space, we can look for different types of errors and diverge into different Prompts depending on the exact error should the student fail to get the answer correct after the Hint state.
@@
 Another thing entirely missing is the “Problematization” strategies described by Reiser (Reiser, 2004) and also used by MathDial.
+While most of these would mainly benefit the cached versions, these can also be made into more specific prompts for the live version.
 Alongside this, the matching requirements for solution alignment can be made stricter to better know where to start applying our intervention.
 This could also come with relaxing the format requirement of “one step per line” on the student solution.
-Both these would shorten the average lengths of the conversation to balance more steps in the strategy space.
+Further, it would be interesting to add in methods to detect when the student ends up completing multiple steps at once.
+All these would shorten the average lengths of the conversation to balance more steps in the strategy space.
 Alternatively, matching can be done using smaller specialized LLMs trained for this task, and this would also apply to correct answer detection and path matching
@@
 In this paper, we propose using LLMs to take over the authoring of the state space for an AutoTutor-like model to minimize human time requirements and build a sample system to showcase our proposal.
-We find that our model performs somewhat better than a Single Prompt GPT4 model in terms of automatic metrics while maintaining full interpretability and modularity in its state space which GPT4 lacks.
+We find that our model performs better than a Single Prompt GPT4 model in terms of automatic metrics while maintaining full interpretability and modularity in its state space which GPT4 lacks.
 Our model also holds up to the desired qualities in its utterances in human evaluation, showing there is high potential in this direction.
@@
 Acknowledgements.
-This work was funded by the Swiss National Science Foundation (Project No.
-197155), and a Responsible AI grant by the Haslerstiftung.
-We extend our gratitude to Jakub Macina from the ETH AI Center, Nico Daheim from UKP Lab, TU Darmstadt, and Farquet Grégoire from Department of Architecture, ETH Zürich for their invaluable feedback.References
+This work was partly funded by the ETH-EPFL Joint Doctoral Program in the Learning Sciences, a grant from the Swiss National Science Foundation (Project No.
+197155) and a Responsible AI grant by the Haslerstiftung.
+We extend our gratitude to Jakub Macina from the ETH AI Center, Nico Daheim from UKP Lab, TU Darmstadt, and Grégoire Farquet from Department of Architecture, ETH Zürich for their invaluable feedback.References
 - (1)
@@
 https://dergipark.org.tr/en/pub/jai/issue/77844/1337500
-- Bloom (1984) Benjamin S Bloom. 1984. The 2 sigma problem: The search for methods of group instruction as effective as one-to-one tutoring. Educational researcher 13, 6 (1984), 4–16.
+- Bloom (1984) Benjamin S Bloom.
+1984.
+The 2 sigma problem: The search for methods of group instruction as effective as one-to-one tutoring.
+Educational researcher 13, 6 (1984), 4–16.
+https://journals.sagepub.com/doi/abs/10.3102/0013189X013006004
 - Bommasani et al.
@@
 https://crfm.stanford.edu/assets/report.pdf
+- Bryson and Hand (2007) Colin Bryson and Len Hand.
+2007.
+The role of engagement in inspiring teaching and learning.
+Innovations in Education and Teaching International 44, 4 (2007), 349–362.
+https://doi.org/10.1080/14703290701602748 arXiv:https://doi.org/10.1080/14703290701602748
 - Cai et al.
@@
 Springer International Publishing, Cham, 593–603.
+https://link.springer.com/chapter/10.1007/978-3-030-22341-0_46
 - Chen et al.
@@
 In Artificial Intelligence in Education.
-Posters and Late Breaking Results, Workshops and Tutorials, Industry and Innovation Tracks, Practitioners, Doctoral Consortium and Blue Sky - 24th International Conference, AIED 2023, Tokyo, Japan, July 3-7, 2023, Proceedings (Communications in Computer and Information Science, Vol.
+Posters and Late Breaking Results Workshops and Tutorials, Industry and Innovation Tracks, Practitioners, Doctoral Consortium and Blue Sky - 24th International Conference, AIED 2023, Tokyo, Japan, July 3-7, 2023, Proceedings (Communications in Computer and Information Science, Vol.
 1831), Ning Wang, Genaro Rebolledo-Mendez, Vania Dimitrova, Noboru Matsuda, and Olga C.
@@
 arXiv preprint arXiv:2110.14168 (2021).
+https://arxiv.org/abs/2110.14168
 - Cui et al.
@@
 arXiv:2309.11495 [cs.CL]
-- Graesser et al. (2000) Arthur Graesser, Natalie Person, and Derek Harter. 2000. Teaching Tactics in AutoTutor. (01 2000).
+- Graesser et al. (2000) Arthur Graesser, Natalie Person, and Derek Harter. 2000. Teaching Tactics in AutoTutor. (01 2000). http://users.sussex.ac.uk/~bend/its2000/graesser.pdf
 - Graesser (2016) Arthur C.
@@
 Springer Berlin Heidelberg, Berlin, Heidelberg, 162–174.
+https://link.springer.com/chapter/10.1007/978-3-540-30139-4_16
 - Kojima et al.
@@
 https://openreview.net/forum?id=fyza2OQ9NI
+- Martin and Bolliger (2018) Florence Martin and Doris Bolliger.
+2018.
+Engagement Matters: Student Perceptions on the Importance of Engagement Strategies in the Online Learning Environment.
+22 (03 2018), 205–222.
+https://doi.org/10.24059/olj.v22i1.1092
 - Nye et al.
@@
 https://doi.org/10.1109/ACCESS.2020.3021383
+- Parsons and Taylor (2011) Jim Parsons and Leah Taylor. 2011. Improving Student Engagement. Current Issues in Education 14, 1 (May 2011). https://cie.asu.edu/ojs/index.php/cieatasu/article/view/745
 - Phung et al.
@@
 IEEE, 181–190.
+https://ieeexplore.ieee.org/abstract/document/7017768
 - Sonkar et al.
@@
 - Zhang et al. (2020) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q. Weinberger, and Yoav Artzi. 2020. BERTScore: Evaluating Text Generation with BERT. arXiv:1904.09675 [cs.CL]
+Appendix A Instructions for Reproduction
+In this section, we list all prompts used by us for various models involving GPT4.
+Parts of the prompt are context specific.
+we mark them using fourth braces eg ¡problem¿.
+We begin with a description of all context specific elements:
+¡problem¿ is replaced by the MWP being solved, eg.
+“Will made 104 dollars mowing lawns over the summer.
+If he spent 41 dollars buying new mower blades, how many 9 dollar games could he buy with the money he had left?”
+¡answer¿ is only the final numerical answer of the current problem. eg “7”
+¡step¿ is the current step of the Solution step space. eg “Divide the money left by the cost of one game to find out how many games he can buy. 63 / 9 = 7 games.”
+¡RHS¿ is the RHS of the equation of the current step
+¡int_sol¿ is the list of all previous steps from the solution step space that have been completed, in each line.
+¡context¿ is the last utterances in the conversation, 2 each from the tutor and the student.
+A.1. GPT4 single prompt
+We use the following system prompt
+Subsequent responses from the student are fed in as user messages, while teacher responses are fed in as assistant messages.
+Conversation stops when the token ”goodbye” is observed in the assistant message.
+A.2. MWPTutor
+The hint, prompt, and final answer for each step, are generated through a sequence of instructions. We start off with the System instruction:
+We resample until the hint checks are satisfied. Thereafter, we perform a correctness check with the instruction:
+If this reproduces the RHS, we store the previous response as the Hint and this response as the Assertion. Finally, we proceed to generate the prompt:
+Again we repeat until the prompt checks are satisfied.
+A.3. MWPTutor
+We generate both prompts and hints using system instructions only. The instruction for prompts is
+Hereafter, for the hint, we say:
+In both cases, we resample until checks are fulfilled
 | Hint | Recent history | Question | Solution Step | Rating |
@@
 6 rows * 6 columns = >36 slots.” What do you think should be the next step?
-Wendy wants to place 20 more than double the number of books in a shelving system with 6 rows and 6 columns. How many books will she need to carry to complete her task? |
-Double the number of slots to find the base number of books before adding the 20 extra books Wendy wants to place. 36 slots * 2 = <<36*2=72>>72 books. |
+Student: The next step would be to represent the number of books Wendy needs to carry as “x”. Then, we can set up an equation to represent the given information. 2x + 20 = 36. |
+Wendy wants to place 20 more than double the number of books in a shelving system with 6 rows and 6 columns.
+How many books will she need to carry to complete her task?
+| Double the number of slots to find the base number of books before adding the 20 extra books Wendy wants to place.
+36 slots * 2 = <<36*2=72>>72 books.
+|
 Grammatical:
@@
 Tutor: 200000 is correct, well done. What is the next step?
-Josh decides to try flipping a house. He buys a house for $80,000 and then puts in $50,000 in repairs. This increased the value of the house by 150%. How much profit did he make? |
-Find out the profit by subtracting the total cost from the new value of the house. $200,000-$130,000 = <<200000-130000=70000>>$70,000. |
+Student: The next step would be to subtract the total cost from the updated value to find the profit made by Josh. |
+Josh decides to try flipping a house.
+He buys a house for $80,000 and then puts in $50,000 in repairs.
+This increased the value of the house by 150%.
+How much profit did he make?
+| Find out the profit by subtracting the total cost from the new value of the house.
+$200,000-$130,000 = <<200000-130000=70000>>$70,000.
+|
 Grammatical:
@@
 Tutor: 65 is correct, well done. What is the next step?
+Student: We need to multiply the number of books sold in the second year by the price of each book, which is $20. |
 Elise has been selling her Dad’s collection of 250 books for three years.
@@
 What’s the total amount of money she earned in the second year?
+| Compute the total amount of money earned in the second year by multiplying the number of books sold by the price per book: <<65 * 20 = 1300>>1300 dollars.
 |
-Compute the total amount of money earned in the second year by multiplying the number of books sold by the price per book: <<65 * 20 = 1300>>1300 dollars. |
 Grammatical:
@@
 Tutor:What should be the first step?
+Student:The first step is to subtract $5 for the arcade tokens from the $90 budget. |
 Morgan’s dad said that she had $90 budgeted for her birthday party.
@@
 How many friends can she invite?
+| Calculate the total cost per person for mini-golf, arcade tokens, and two go-kart rides.
+5 + 5 + (10 x 2) = <<5 + 5 + (10 * 2)=30>>30 dollars per person.
 |
-Calculate the total cost per person for mini-golf, arcade tokens, and two go-kart rides. 5 + 5 + (10 x 2) = <<5 + 5 + (10 * 2)=30>>30 dollars per person. |
 Grammatical:
 YYY Relevant: YYY Not redundant: YYY Not revealing: NYN |
+| Hint | Recent history | Question | Solution Step | Rating |
+| Exactly, let’s keep the momentum going.
+Now, think about how we could figure out the total number of books Elise managed to sell over all three years.
+What should we do with the number of unsold books to find that out?
+|
+Tutor: 90 is correct, well done. What is the next step?
+Student: The next step is to calculate how many books were sold in the second and third year combined. |
+Elise has been selling her Dad’s collection of 250 books for three years.
+Each book sells at 20$, and she sold twice as many books in the first year as she has sold in the current year.
+There are currently 50 unsold books, and her sales number this year is 45.
+What’s the total amount of money she earned in the second year?
+| Determine the total number of books sold over three years by subtracting unsold books from the initial collection: 200 books.
+|
+Grammatical:
+NNY Relevant: YYY Not redundant: YYY Not revealing: NYY |
+| How about we start by figuring out how many hotdogs were prepared for half of the guests to have two? Remember, the caterer was initially expecting 36 guests. |
+Tutor: What should be the first step?
+Student: The first step should be to figure out how many hotdogs the caterer prepared in total. |
+A food caterer was told to prepare gourmet hot dogs for 36 guests.
+While most people would only eat one hotdog, he prepared enough for half of the guests to be able to have two hotdogs.
+However, 40 guests showed up, and everyone wanted a second hotdog.
+How many guests did not get a second hotdog?
+| Calculate the total number of hotdogs the caterer prepared for half the guests to have two.
+Total hotdogs prepared = Number of guests initially expected + Extra hotdogs for half of the guests = 36 + (36/2) = 54 hotdogs.
+|
+Grammatical:
+YYY Relevant: YNN Not redundant: YYY Not revealing: YYY |
+| Actually, before we look at how many Post-it notes she used at work, can you think of how we might find out how many Post-it notes Candice had after she stopped at the store but before she started using them at work?
+Remember, she started with a certain number and then added more from the package she bought.
+|
+Tutor:What should be the first step?
+Student:The first step should be to figure out how many post-it notes Candice used at work. |
+Candice put 80 post-it notes in her purse before she headed out to her job at the coffee shop.
+On her way, she stopped off at the store and purchased a package of Post-it notes; At work, she placed a single Post-it note on each of 220 different cups of coffee.
+If she had 23 post-it notes remaining overall, how many Post-it notes were in the package that she purchased?
+| Determine the total number of Post-it notes used at the coffee shop by subtracting the remaining Post-it notes from the initial amount Candice had.
+80 initial Post-it notes minus 23 remaining equals 57 Post-it notes used.
+|
+Grammatical:
+NYY Relevant: YYY Not redundant: NNN Not revealing: YYY |
+| That’s a great thought process!
+Now, can you figure out the number of books Elise sold in the second year?
+Remember, you’ll need to use the total sales over three years and subtract the sales from the first and third years from this total.
+|
+Tutor: 200 is correct, well done. What is the next step?
+Student: The next step is to figure out how many books were sold in the second year by subtracting the number of books sold in the first and third year from the total number of books sold over all three years.
+|
+Elise has been selling her Dad’s collection of 250 books for three years.
+Each book sells at 20$, and she sold twice as many books in the first year as she has sold in the current year.
+There are currently 50 unsold books, and her sales number this year is 45.
+What’s the total amount of money she earned in the second year?
+| Calculate how many books were sold in the second year by subtracting the first and third year sales from the total sales: 65 books..
+|
+Grammatical:
+YYY Relevant: YYY Not redundant: YYY Not revealing: NNN |
 | Hint | Recent history | Question | Solution Step | Rating |
@@
 Tutor: Tell me this instead: What is the total monthly expense for Madeline’s dog when combining the costs of food, treats, and medicine?
+Student: Oh, right. So it’s $25 + $20 + $100 = $145 per month. |
 Every month, Madeline has to buy food, treats, and medicine for her dog.
@@
 How much money does Madeline spend on her dog per year if there are 4 weeks in a month?
-|
 | Add the monthly costs of food, treats, and medicine to find the total monthly expenses for Madeline’s dog.
@@
 Tutor: Tell me this instead: What is the number of games Bobby has if you take away 5 from three times the number of Brian’s games?
+Student: If we take away 5 from three times the number of Brian’s games, we get 3*20-5 = 55 video games for Bobby. |
 Brian’s friend Bobby has 5 fewer than 3 times as many video games as Brian does. If Brian has 20 video games but lost 5 right before the comparison was made, how many does Bobby have? |
 Subtract 5 from the tripled number of Brian’s games to find out how many games Bobby has.
+45-5=<<45-5=40>> 40 video games. |
 Grammatical:
 YYY Specific: YYY Helpful: YYY Not revealing: NNY |
-An example solution and utterance space from the MultiArith dataset obtained using GPT4
+| Hint | Recent history | Question | Solution Step | Rating |
+| It seems there’s a little confusion.
+To clarify, since the two pairs of heels together cost five dollars less than the boots, to find the cost of the boots, we actually need to add $5 to the total cost of the heels, not subtract it.
+Given this, what operation should we perform with the $99 to find the cost of the boots?
+|
+Tutor: Actually, remember the problem stated that the two pairs of high heels together cost five dollars less than the boots.
+With that in mind, what do you think we should do with the $99, which is the total cost of the two pairs of heels, to find out how much more expensive the boots are compared to this total?
+Student: Oh, I see. We should subtract $5 from $99 to find the cost of just the boots. So the final answer is $94. |
+Gloria is shoe shopping when she comes across a pair of boots that fit her shoe budget.
+However, she has to choose between the boots and two pairs of high heels that together cost five dollars less than the boots.
+If one pair of heels costs $33 and the other costs twice as much, how many dollars are the boots?
+| Determine the cost of the boots by adding five dollars to the total cost of the two pairs of heels.
+99 + 5 = 104 dollars |
+Grammatical:
+NNY Specific: YYY Helpful: YYY Not revealing: YNY |
+| Actually, there seems to be a misunderstanding about how to double the slots.
+To double the number of slots doesn’t involve dividing them by 2.
+Instead, think about what ”double” means in terms of multiplication.
+Can you try recalculating how many books would fit if we double the space available, before adding the extra 20 books Wendy wants?
+|
+Tutor: That’s a good attempt, but let’s consider Wendy’s plan a little more closely.
+She wants to place 20 more than double the number of books based on the slots available.
+Given that we’ve calculated the total slots, how do you think we should figure out what double the number of slots would represent in terms of books?
+Student:Oh, I see.
+We would need to divide the total slots by 2 to get the number of books that would fit in double the slots.
+So, it would be 36/2 = 18 books.
+Then, we add 20 to get the final number of books needed, which would be 18 + 20 = 38 books.
+|
+Wendy wants to place 20 more than double the number of books in a shelving system with 6 rows and 6 columns.
+How many books will she need to carry to complete her task?
+| Double the number of slots to find the base number of books before adding the 20 extra books Wendy wants to place.
+36 slots * 2 = 72 books.
+|
+Grammatical:
+YYY Specific: NYN Helpful: NYY Not revealing: YNY |
+| Actually, it seems there’s a misunderstanding in how we’re adding up the bids.
+You’ve done well to identify the amounts for each of Carmen’s subsequent bids, but we also need to consider the total cost by adding all of her bids together, not just the final bid amount.
+Remember, the question asks for the total amount Carmen pays, which includes her initial bid of $200 plus the additional bids she places.
+Can you calculate the sum of all her bids to find the total cost?
+|
+Tutor: Actually, it seems there’s a misunderstanding in how we’re adding up the bids.
+You’ve done well to identify the amounts for each of Carmen’s subsequent bids, but we also need to consider the total cost by adding all of her bids together, not just the final bid amount.
+Remember, the question asks for the total amount Carmen pays, which includes her initial bid of $200 plus the additional bids she places.
+Can you calculate the sum of all her bids to find the total cost?
+Student: Yes, so the second bid is $250, the third bid is $300, and the fourth bid is $350. So the total cost for Carmen to win the desk is $350. |
+Carmen goes to an auction to win an antique desk.
+She accepts the opening bid of $200 and continues bidding until she wins.
+The bids on the desk rise by $50 each time and 3 other people each bid once.
+Carmen bids after each of the 3 other people and eventually wins.
+How much money, in dollars, does the desk cost her?
+| Carmen accepts the opening bid, which is 200 |
+Grammatical:
+YYY Specific: YYY Helpful: YNN Not revealing: YYY |
+| It looks like there might have been a small mistake in your calculation.
+Remember to first convert the annual interest rate to a monthly rate by dividing it by 12, and then multiply by the number of months, which is 5 in this case, and finally apply it to the borrowed amount to find the total interest.
+Could you try recalculating the total interest for the five-month period using these steps?
+|
+Tutor: Yes, you’re on the right track!
+Now, with that approach, calculate the total interest for the five-month period using the annual interest rate of 10%.
+Remember, after finding the monthly rate, multiply it by the number of months to get the total interest.
+What do you get?
+Student: I got $152.08 for the total interest. |
+Karan borrowed $3,650 for five months at an interest rate of 10%. She has to pay an equal amount every month. How much does she have to pay per month? |
+Calculate the total interest for the 5-month period: 3650 * 10% = 365 dollars.
+45-5=<<45-5=40>> 40 video games. |
+Grammatical:
+YYY Specific: YYY Helpful: YYY Not revealing: NNY |
+[breaklines, fontsize=]json ”question”: ” Will made 104 dollars mowing lawns over the summer.
+If he spent 41 dollars buying new mower blades, how many 9-dollar games could he buy with the money he had left?
+”, ”answer”: 7, ”solution”: ”step”: [ ”Calculate the money left after Will spent on the mower blades by subtracting the cost of the blades from the initial amount he made.
+104 - 41 = ¡¡104 - 41=63¿¿63 dollars left.” ], ”follow-ups”: [ ”step”: [ ”Divide the money left by the cost of one game to find out how many games he can buy.
+63 / 9 = ¡¡63 / 9=7¿¿7 games.
+QED” ], ”follow-ups”: [], ”hint”: ”How many games can he buy with the money he has left if one game costs nine dollars?”, ”prompt”: ”To find out how many games he can purchase, determine the number of times the cost of one game goes into the total amount of money he has left.”, ”assertion”: ”He can buy 7 games with the money left if one game costs nine dollars.” , ”step”: [ ”Determine how many 9-dollar games Will can buy by dividing the money he has left by the cost of one game: 63 ÷ 9 = ¡¡63 ÷ 9=7¿¿7 games.
+QED” ], ”follow-ups”: [], ”hint”: ”How many 9 dollar games can Will buy with the money he has left?”, ”prompt”: ”Consider how many times the cost of one game goes into the total amount of money Will has to spend.”, ”assertion”: ”Will can buy 7 nine-dollar games with the money he has left.” ], ”hint”: ”How much money is left after Will spends on mower blades if he starts with 104 dollars?”, ”prompt”: ”Consider the initial amount Will had and subtract the cost of the mower blades he purchased to find out the remaining amount.”, ”assertion”: ”Will has 63 dollars left after spending on the mower blades.” An example solution and utterance space from the MultiArith dataset obtained using GPT4